### v1

# Direct reconstruction of anisotropic self-adjoint inclusions in the Calderón problem

Abstract.
We extend the monotonicity method for direct exact reconstruction of inclusions in the partial data Calderón problem, to the case of anisotropic conductivities in any spatial dimension $d\geq 2$. Specifically, from a local Neumann-to-Dirichlet map, we give reconstruction methods of inclusions based on unknown anisotropic self-adjoint perturbations to a known anisotropic conductivity coefficient. A main assumption is a definiteness condition for the perturbations near the outer inclusion boundaries. This additionally provides new insights into the non-uniqueness issues of the anisotropic Calderón problem.
Keywords: Anisotropic Calderón problem, inclusion detection, monotonicity method.
2020 Mathematics Subject Classification: 35R30, 35R05, 47H05.
1. Introduction
The anisotropic Calderón problem, the inverse problem of determining an elliptic matrix-valued conductivity coefficient $A$ from boundary measurements of the conductivity equation
[…]
is of major interest in the inverse problems community. Unlike the isotropic (scalar-valued) case, the anisotropic Calderón problem suffers from significant non-uniqueness issues. Indeed if boundary measurements are taken on $\Gamma\subseteq\partial\Omega$, for any $H^{1}$-diffeomorphism $\Phi\colon\overline{\Omega}\to\overline{\Omega}$ for which $\Phi|_{\Gamma}$ is the identity, the conductivity coefficient defined via the push-forward
$\widetilde{A}=\frac{\mathop{\textup{D}\Phi}\nolimits A\mathop{\textup{D}\Phi}\nolimits^{\textup{T}}}{\lvert\det\mathop{\textup{D}\Phi}\nolimits\rvert}\circ\Phi^{-1}$ (1.1)
produces the same local Neumann-to-Dirichlet (ND) map as the coefficient $A$. For $\Gamma=\partial\Omega$ and with spatial dimension $d=2$, the obstructions to uniqueness are characterized by (1.1) for real-valued symmetric $L^{\infty}$-conductivities [4, 37]. In spatial dimension $d\geq 3$, the full extent of the non-uniqueness issues (beyond real-analytic coefficients) remains one of the most important unanswered questions in the inverse problems community [32, 39]. In a fractional Calderón problem, using the stronger non-local properties, the obstructions to uniqueness are characterized for $d\geq 2$ [9]. Despite knowledge on the possible obstructions, the non-uniqueness has prevented development of practically useful reconstruction methods. Unique recovery is only guaranteed with very strong restrictions on the considered class of coefficients [2, 3, 10].
In this paper we study the exact reconstruction of anisotropic inclusions in any spatial dimension $d\geq 2$. We reconstruct the outer shape $D^{\bullet}$ of
$D=\mathop{\textup{supp}}(A_{D}-A_{0}),$
from knowledge of $A_{0}$ and the local ND map $\Lambda(A_{D})$ associated with the unknown conductivity $A_{D}$. Here $D^{\bullet}$ is the smallest closed set with connected complement such that $D\subseteq D^{\bullet}$. We give direct reconstruction methods (Theorems 3.6 and 3.7) based on the monotonicity of ND maps. The main assumptions are self-adjointness of the coefficients, that $A_{0}$ satisfies a unique continuation principle, and that $A_{D}-A_{0}$ satisfies a local uniform definiteness condition inside $D$ near $\partial D^{\bullet}$; see Assumption 3.4 for the precise statement. In particular, inside $D$ we do not assume more regularity than $L^{\infty}$ for $A_{D}$, and $A_{D}-A_{0}$ can even be indefinite at a positive distance away from $\partial D^{\bullet}$.
As an indirect consequence: if such a perturbation $A_{D}-A_{0}$, to a given “background conductivity” $A_{0}$, is known to satisfy the definiteness condition near $\partial D^{\bullet}$, this is enough to ensure that all coefficient-transformations (1.1), compatible with this condition, satisfy
$\mathop{\textup{supp}}(\widetilde{A}_{D}-A_{0})^{\bullet}=D^{\bullet}.$
Moreover, the inclusion detection problem is also unaffected by any other potential forms of non-uniqueness that could arise from the anisotropic Calderón problem when $d\geq 3$. Hence, for inclusion detection, practical reconstruction is sensible in a very high generality compared to reconstructing an anisotropic conductivity.
To give some intuition on why this is the case, let
$\widetilde{D}=\Phi(D)$
be the transformed inclusion under $\Phi$. If $\widetilde{D}^{\bullet}\neq D^{\bullet}$ then $\Phi$ will fail to be the identity in an open set intersecting $\partial D^{\bullet}$ (the part where it is deformed by $\Phi$), which also changes the background conductivity in $\widetilde{A}_{D}$ in that part of the domain via (1.1). From our results we conclude in this situation, that for
$\widehat{D}=\mathop{\textup{supp}}(\widetilde{A}_{D}-A_{0}),$
if $\widehat{D}^{\bullet}\neq D^{\bullet}$, the definiteness condition cannot be satisfied inside $\widehat{D}$ near $\partial\widehat{D}^{\bullet}$. In particular, it may fail in part of the domain where the background conductivity in $\widetilde{A}_{D}$ is modified.
In the context of Helmholtz scattering, see also the uniqueness results in [21, 6] for inclusion detection with some anisotropic coefficients.
1.1. The monotonicity method and related results
The reconstruction methods we present are based on the monotonicity of ND maps via the following operator inequalities (Proposition 5.5):
$\int_{\Omega}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x\leq\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle\leq\int_{\Omega}A_{2}A_{1}^{-1}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x,$
where $\Lambda_{j}=\Lambda(A_{j})$ and $u_{2}$ is the electric potential for conductivity $A_{2}$ and Neumann condition $f$. The definiteness properties of $A_{2}A_{1}^{-1}(A_{2}-A_{1})$ turn out to be the same as for $A_{2}-A_{1}$ (Proposition 5.6). The main idea in the outer approach of the monotonicity method, given in our main results Theorems 3.6 and 3.7, is to construct test-operators that use the monotonicity to check if a given test-set $C$ fully contains $D^{\bullet}$ or not. This leads to an if and only if statement related to operator inequalities with the test-operators and the ND map for the unknown perturbed conductivity. In Theorem 3.7 the test-operators are linearized, which is beneficial for numerical implementation.
The proofs rely on the theory of localized potentials [20], for the existence of certain localizing solutions to the conductivity equation, based on a unique continuation principle. We generalize the result on localized potentials to anisotropic (even non-self-adjoint) coefficients in Theorem 6.3.
Under much stronger assumptions, we also give an inner approach of the monotonicity method in Theorems 4.2 and  4.3, which checks if a given open set $B$ is contained in $D^{\bullet}$.
The monotonicity method is well-established for isotropic conductivities, and for other similar inverse coefficient problems. We mention some of the results here, with a focus on the Calderón problem. In [38] they use the monotonicity to give bounds on inclusions, and later [26, 29] prove that it is an exact reconstruction method for finite perturbations to the background conductivity, using the theory of localized potentials. It also includes the case with both positive and negative perturbations, if lower and upper bounds are known for the perturbed conductivity. This is expanded upon in [7], and now the perturbed conductivity can simultaneously have parts with finite positive and negative perturbations, as well as extreme parts that are perfectly insulating and perfectly conducting. The need for lower and upper bounds is also removed. Finally, in [18] the method is generalized to reconstruct collections of Lipschitz cracks.
The method has also been used in other specialized directions: degenerate and singular perturbations based on $A_{2}$-Muckenhoupt weights [15], inclusions in a fractional Calderón problem [27, 28], inclusions in a $p$-Laplace equation [5], for nonlinear materials [36], and for local uniqueness in an inverse coefficient problem for a non-resonant Schrödinger equation [25]. There are also rigorous connections to practically relevant electrode models [16, 17, 22, 30].
Certain reconstruction methods for the partial data isotropic Calderón problem also rely on the same type of methodology. For piecewise constant layered conductivities, there is an exact reconstruction method for both conductivity coefficient and the piecewise constant partition [11, 12]. In a finite-dimensional settting (where the partition is known), a general piecewise constant conductivity is reconstructed in [13], and likewise via a reformulation to a convex semidefinite optimization problem in [23].
1.2. Remarks on notation
All considered vector spaces are complex; this includes the classes of conductivity coefficients. At all occurrences of using support/supremum/infimum, we make use of the essential support/supremum/infimum.
We use the convention that inner products on complex Hilbert spaces are linear in the first entry and anti-linear in the second. We denote the Euclidean inner product as $z_{1}\cdot\overline{z_{2}}$ for $z_{1},z_{2}\in\mathbb{C}^{d}$, in particular the “dot” is bilinear. The Euclidean norm is denoted $\lvert\,\cdot\,\rvert$.
For self-adjoint operators $A,B\in\mathscr{L}(H)$ for a Hilbert space $H$, we write $A\geq B$ when ABA-B is positive semidefinite; this is the Loewner ordering of such operators. Moreover, for $A$ being a positive definite operator, $A^{1/2}$ denotes its unique positive definite square root.
2. The forward problem
Let $\Omega$ be a bounded Lipschitz domain in $\mathbb{R}^{d}$, for $d\in\mathbb{N}\setminus\{1\}$, with connected complement. For a bounded matrix-valued function $A\in L^{\infty}(\Omega)^{d\times d}$, we use the norm
$\lVert A\rVert_{*}=\sup_{x\in\Omega}\lVert A(x)\rVert_{2},$
where $\lVert\,\cdot\,\rVert_{2}$ is the usual spectral norm for matrices, i.e. the operator norm on Euclidean space. Hence, we have
$\lvert A(x)\xi\rvert\leq\lVert A\rVert_{*}\lvert\xi\rvert$
for all $\xi\in\mathbb{C}^{d}$ and a.e. $x\in\Omega$.
We use the notation
$A^{\textup{R}}=\frac{1}{2}\bigl{(}A+A^{*}\bigr{)}\quad\text{and}\quad A^{\textup{I}}=\frac{1}{2\mathrm{i}}\bigl{(}A-A^{*}\bigr{)},$
such that
$A=A^{\textup{R}}+\mathrm{i}A^{\textup{I}},$
where we note that $A^{\textup{R}}$ and $A^{\textup{I}}$ are self-adjoint as matrices, i.e. Hermitian at every point in $\Omega$. This gives rise to the following general class of anisotropic coefficients:
$\mathcal{H}(\Omega)=\{\,A\in L^{\infty}(\Omega)^{d\times d}\mid\exists c>0\colon A^{\textup{R}}\geq cI\text{ in the Loewner order in $\Omega$}\,\}.$
The condition on $A$, the uniform positive definiteness of $A^{\textup{R}}$, can also explicitly be written:
$\inf_{x\in\Omega}\mathop{\textup{Re}}\bigl{(}A(x)\xi\cdot\overline{\xi}\bigr{)}=\inf_{x\in\Omega}A^{\textup{R}}(x)\xi\cdot\overline{\xi}\geq c\lvert\xi\rvert^{2}$
for all $\xi\in\mathbb{C}^{d}$; there are other classes of strongly elliptic coefficients (when restricting to $\xi\in\mathbb{R}^{d}$) that we do not consider here, see e.g. [34, chapter 4]. We also have the subset where $A$ is self-adjoint, i.e. when $A=A^{\textup{R}}$:
$\mathcal{H}_{\textup{SA}}(\Omega)=\{\,A\in\mathcal{H}(\Omega)\mid A^{\textup{I}}=0\,\}.$
Consider the partial data anisotropic conductivity problem for $A\in\mathcal{H}(\Omega)$:
[…]
Here $\nu$ is the outer unit normal to $\Omega$, $\Gamma\subseteq\partial\Omega$ is a non-empty open boundary piece, and the mean-free current density $f$ belongs to
$L^{2}_{\diamond}(\Gamma)=\{\,f\in L^{2}(\Gamma)\mid\langle f,1\rangle=0\,\}.$
We denote by $\langle\,\cdot\,$$\cdot\,\rangle$ the standard inner product on $L^{2}(\Gamma)$, and $\lVert\,\cdot\,\rVert$ is the associated norm. We likewise denote
$H^{1}_{\diamond}(\Omega)=\{\,u\in H^{1}(\Omega)\mid\langle u|_{\Gamma},1\rangle=0\,\}.$
A Poincaré inequality, related to the mean-free condition in $H^{1}_{\diamond}(\Omega)$, enables that
$\lVert v\rVert_{H^{1}_{\diamond}(\Omega)}=\Bigl{(}\int_{\Omega}\lvert\nabla v\rvert^{2}\,\mathrm{d}x\Bigr{)}^{1/2}$
becomes a norm on $H^{1}_{\diamond}(\Omega)$, equivalent with the usual $H^{1}$-norm. Moreover, the following continuity and coercivity estimates hold for all $w,v\in H^{1}_{\diamond}(\Omega)$:
$\bigl{\lvert}\int_{\Omega}A\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x\bigr{\rvert}$ $\leq\lVert A\rVert_{*}\lVert w\rVert_{H^{1}_{\diamond}(\Omega)}\lVert v\rVert_{H^{1}_{\diamond}(\Omega)},$
$\bigl{\lvert}\int_{\Omega}A\nabla v\cdot\overline{\nabla v}\,\mathrm{d}x\bigr{\rvert}$ $\geq\mathop{\textup{Re}}\int_{\Omega}A\nabla v\cdot\overline{\nabla v}\,\mathrm{d}x\geq c\lVert v\rVert_{H^{1}_{\diamond}(\Omega)}^{2}.$
By the Lax–Milgram lemma, there exists a unique solution $u=u_{f}^{A}\in H^{1}_{\diamond}(\Omega)$ to the corresponding weak problem associated to (2.1):
$\int_{\Omega}A\nabla u\cdot\overline{\nabla v}\,\mathrm{d}x=\langle f,v\rangle$ (2.2)
for all $v\in H^{1}_{\diamond}(\Omega)$. We let $\Lambda(A)\in\mathscr{L}(L^{2}_{\diamond}(\Gamma))$ denote the associated local ND map,
$\Lambda(A)f=u_{f}^{A}|_{\Gamma},$
and the forward problem is the nonlinear mapping $\Lambda\colon\mathcal{H}(\Omega)\to\mathscr{L}(L^{2}_{\diamond}(\Gamma))$. From (2.2) it holds that for any $A_{1},A_{2}\in\mathcal{H}(\Omega)$ and $f,g\in L^{2}_{\diamond}(\Gamma)$, we have
$\langle f,\Lambda(A_{1})g\rangle=\int_{\Omega}A_{2}\nabla u_{f}^{A_{2}}\cdot\overline{\nabla u_{g}^{A_{1}}}\,\mathrm{d}x.$ (2.3)
Thus the adjoint operator satisfies
$\langle f,\Lambda(A_{1})^{*}g\rangle=\overline{\langle g,\Lambda(A_{1})f\rangle}=\int_{\Omega}\overline{A_{2}\nabla u_{g}^{A_{2}}}\cdot\nabla u_{f}^{A_{1}}\,\mathrm{d}x.$
Of particular interest are the quadratic forms
$\langle f,\Lambda(A)f\rangle=\int_{\Omega}A\nabla u_{f}^{A}\cdot\overline{\nabla u_{f}^{A}}\,\mathrm{d}x\quad\text{and}\quad\langle f,\Lambda(A)^{*}f\rangle=\int_{\Omega}A^{*}\nabla u_{f}^{A}\cdot\overline{\nabla u_{f}^{A}}\,\mathrm{d}x.$
Specifically,
$\Lambda(A)^{*}=\Lambda(A^{*}),$
meaning that $\Lambda(A)$ is self-adjoint if $A\in\mathcal{H}_{\textup{SA}}(\Omega)$. We also write
$\Lambda^{\!\textup{R}}(A)=\frac{1}{2}\bigl{(}\Lambda(A)+\Lambda(A)^{*}\bigr{)}\quad\text{and}\quad\Lambda^{\!\textup{I}}(A)=\frac{1}{2\mathrm{i}}\bigl{(}\Lambda(A)-\Lambda(A)^{*}\bigr{)},$
so for the forward problem we have
$\Lambda=\Lambda^{\!\textup{R}}+\mathrm{i}\Lambda^{\!\textup{I}},$
where $\Lambda^{\!\textup{R}}(A)$ and $\Lambda^{\!\textup{I}}(A)$ are self-adjoint operators. In terms of the quadratic forms we thus have
$\langle f,\Lambda^{\!\textup{R}}(A)f\rangle$ $=\mathop{\textup{Re}}\langle f,\Lambda(A)f\rangle=\int_{\Omega}A^{\textup{R}}\nabla u_{f}^{A}\cdot\overline{\nabla u_{f}^{A}}\,\mathrm{d}x=\int_{\Omega}\lvert(A^{\textup{R}})^{1/2}\nabla u_{f}^{A}\rvert^{2}\,\mathrm{d}x,$ (2.4)
$\langle f,\Lambda^{\!\textup{I}}(A)f\rangle$ $=-\mathop{\textup{Im}}\langle f,\Lambda(A)f\rangle=-\int_{\Omega}A^{\textup{I}}\nabla u_{f}^{A}\cdot\overline{\nabla u_{f}^{A}}\,\mathrm{d}x.$ (2.5)
As one would expect, when multiplying with a complex scalar $\kappa\in\mathbb{C}$, we get
$(\kappa\Lambda)^{\textup{R}}=\kappa^{\textup{R}}\Lambda^{\!\textup{R}}-\kappa^{\textup{I}}\Lambda^{\!\textup{I}}\quad\text{and}\quad(\kappa\Lambda)^{\textup{I}}=\kappa^{\textup{I}}\Lambda^{\!\textup{R}}+\kappa^{\textup{R}}\Lambda^{\!\textup{I}}.$
3. Outer reconstruction of general anisotropic inclusions
In the following we will assume that a so-called “background conductivity” $A_{0}\in\mathcal{H}_{\textup{SA}}(\Omega)$ is known. For an unknown $A_{D}\in\mathcal{H}_{\textup{SA}}(\Omega)$, we call
$D=\mathop{\textup{supp}}(A_{D}-A_{0})$
the inclusions, noting that $D$ may have several connected components. Our reconstruction methods will relate to the so-called outer shape of $D$.
Definition 3.1.
The outer shape $D^{\bullet}$ of $D$, is the smallest closed set with connected complement such that $D\subseteq D^{\bullet}$.
One of the key assumptions we will need, required for the existence of certain localized solutions, is that the background conductivity $A_{0}$ satisfies a unique continuation principle.
Definition 3.2.
Suppose $U\subseteq\overline{\Omega}$ is a relatively open and connected set that intersects $\Gamma$. We say that $A\in\mathcal{H}(\Omega)$ satisfies the weak unique continuation principle (UCP) in $U$ for the conductivity equation, provided that only the trivial solution of
[…]
can be identically zero in a non-empty open subset of $U$, and only the trivial solution has vanishing Cauchy data, $u$ and $\nu\cdot(A\nabla u)$, on a non-empty open part of $\partial U\cap\Gamma$. If $U=\overline{\Omega}$ we just say that $A$ satisfies the UCP.
Remark $3.$
In spatial dimension $d=2$, the UCP holds for general real-valued $A\in\mathcal{H}_{\textup{SA}}(\Omega)$ (and even in some non-self-adjoint cases), see [1].
In spatial dimension $d\geq 3$, the UCP holds for real-valued Lipschitz continuous coefficients $A\in\mathcal{H}_{\textup{SA}}(\Omega)$, see [19], and with counterexamples to the UCP for $\alpha$-Hölder continuous coefficients with $\alpha\in(0,1)$, see [35, 33].
We note that there also exist UCP results related to classes of complex coefficients, e.g. with principally normal symbol [31, chapter 28].
We define the class of admissible test-inclusions, used for the reconstruction methods, as
[…]
Assumption 3.4.
  1. (i)
    Assume there are known bounds on $A_{D}$, i.e. scalars $0<\alpha<\beta$ such that
    $\alpha I\leq A_{D}\leq\beta I$
    in the Loewner order in $\Omega$.
  2. (ii)
    Assume that $D\Subset\Omega$ and is the closure of an open set.
  3. (iii)
    Assume for every open neighborhood $W$ of $x\in\partial D^{\bullet}$, there exists a relatively open connected set $V\subset D^{\bullet}\cap W$ that intersects $\partial D^{\bullet}$, satisfying either of two options:
    1. (a)
      $A_{D}-A_{0}$ is positive semidefinite in $V$, and there exists an open ball $B\subset V$ on which $A_{D}-A_{0}$ is uniformly positive definite.
    2. (b)
      $A_{D}-A_{0}$ is negative semidefinite in $V$, and there exists an open ball $B\subset V$ on which $A_{D}-A_{0}$ is uniformly negative definite.
  4. (iv)
    Assume that $A_{0}$ satisfies the UCP.
We give a few remarks on these assumptions.
Remark 3.5 (Remarks on Assumption 3.4).
  1. (a)
    In case $\Gamma=\partial\Omega$, then $C\Subset\Omega$ in the definition of $\mathcal{A}$ can be replaced by $C\subseteq\overline{\Omega}$, and $D\Subset\Omega$ in Assumption 3.4(ii) can be replaced by $D\subseteq\overline{\Omega}$. Also, here “connected complement” is in $\mathbb{R}^{d}$.
  2. (b)
    Note that Assumption 3.4(iii) only relates to what happens close to $\partial D^{\bullet}$, so further into $D$ we may have that $A_{D}-A_{0}$ is e.g. indefinite. Moreover, the definiteness condition is not exactly at $\partial D^{\bullet}$ but just close to $\partial D^{\bullet}$, meaning that there can be a continuous deviation from $A_{0}$ rather than a jump condition; in the isotropic case this mimics the condition of a strict local increase/decrease from the background conductivity when entering $D$ from the outside [15].
For a measurable set $C$, we define the test-coefficients
$A^{-}_{C}=\begin{dcases}\alpha I&\text{in }C\\ A_{0}&\text{in }\Omega\setminus C\end{dcases}\quad\text{and}\quad A^{+}_{C}=\begin{dcases}\beta I&\text{in }C\\ A_{0}&\text{in }\Omega\setminus C,\end{dcases}$
and the corresponding test-operators
[…]
Theorem 3.6.
Under Assumption 3.4(i), for any measurable $C\subseteq\overline{\Omega}$ we have
$D\subseteq C\quad\text{implies}\quad\Lambda^{-}_{C}\geq\Lambda(A_{D})\geq\Lambda^{+}_{C}$.
Under all of Assumption 3.4, for any $C\in\mathcal{A}$ we have
$\Lambda^{-}_{C}\geq\Lambda(A_{D})\geq\Lambda^{+}_{C}\quad\text{implies}\quad D\subseteq C.$
Proof.
The proof is given in Section 7. ∎
We also introduce linearized versions of the test-operators, which are more suitable for numerical computations. Note that for $A\in\mathcal{H}_{\textup{SA}}(\Omega)$ and $B\in L^{\infty}(\Omega)^{d\times d}$, the Frechét derivative of $\Lambda$ evaluated at $A$ and in direction $B$ is given by (c.f. [14]):
$\langle f,\textup{D}\mkern-1.5mu\Lambda(A;B)f\rangle=-\int_{\Omega}B^{*}\nabla u_{f}^{A}\cdot\overline{\nabla u_{f}^{A}}\,\mathrm{d}x.$
To shorten the notation in the following result, we introduce the test-operators
$\textup{D}\mkern-1.5mu\Lambda^{+}_{C}$ $=\textup{D}\mkern-1.5mu\Lambda\bigl{(}A_{0};(\beta I-A_{0})\chi_{C}\bigr{)},$
$\textup{D}\mkern-1.5mu\Lambda^{-}_{C}$ $=\textup{D}\mkern-1.5mu\Lambda\bigl{(}A_{0};(A_{0}-\tfrac{\beta^{2}}{\alpha}I)\chi_{C}\bigr{)},$
where $\chi_{C}$ is a characteristic function on the set $C$.
Theorem 3.7.
Pick $0<\alpha<\beta$ such that Assumption 3.4(i) holds for both $A_{D}$ and $A_{0}$. For any measurable $C\subseteq\overline{\Omega}$ we have
$D\subseteq C\quad\text{implies}\quad\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\geq\Lambda(A_{D})-\Lambda(A_{0})\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}$.
Under all of Assumption 3.4, for any $C\in\mathcal{A}$ we have
$\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\geq\Lambda(A_{D})-\Lambda(A_{0})\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\quad\text{implies}\quad D\subseteq C.$
Proof.
The proof is given in Section 8. ∎
Remark 3.8.
Under the given assumptions, Theorem 3.6 gives
$D^{\bullet}=\cap\,\{\,C\in\mathcal{A}\mid\Lambda^{-}_{C}\geq\Lambda(A_{D})\geq\Lambda^{+}_{C}\,\}$,
while Theorem 3.7 gives
$D^{\bullet}=\cap\,\{\,C\in\mathcal{A}\mid\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\geq\Lambda(A_{D})-\Lambda(A_{0})\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\,\}$.
Moreover, from the proofs we also conclude:
  • If in Assumption 3.4(iii) there is only positive definiteness near $\partial D^{\bullet}$, we only need to check the inequalities $\Lambda(A_{D})\geq\Lambda^{+}_{C}$ or $\Lambda(A_{D})-\Lambda(A_{0})\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}$.
  • If in Assumption 3.4(iii) there is only negative definiteness near $\partial D^{\bullet}$, we only need to check the inequalities $\Lambda^{-}_{C}\geq\Lambda(A_{D})$ or $\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\geq\Lambda(A_{D})-\Lambda(A_{0})$.
4. Inner reconstruction of uniformly definite anisotropic inclusions
Under much stronger definiteness assumptions (not just near $\partial D^{\bullet}$ as before), we can also use the inner approach of the monotonicity method. We focus on the linearized version, so that the UCP requirement is only related to the background conductivity $A_{0}$.
Assumption 4.1.
  1. (i)
    Assume that we have either:
    1. (a)
      Positive definite inclusions, $A_{D}-A_{0}\geq cI$ in the Loewner order in $D$, for known $c>0$.
    2. (b)
      Negative definite inclusions, $A_{D}-A_{0}\leq-cI$ in the Loewner order in $D$, for known $c>0$.
  2. (ii)
    Assume that $D\Subset\Omega$ and is the closure of an open set.
  3. (iii)
    Assume that $A_{0}$ satisfies the UCP.
Again, to shorten the notation in the following results, we introduce the test-operators
$\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}$ $=\textup{D}\mkern-1.5mu\Lambda\bigl{(}A_{0};c(\tfrac{\alpha}{\beta})^{2}\chi_{B}I\bigr{)},$
$\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}$ $=\textup{D}\mkern-1.5mu\Lambda\bigl{(}A_{0};-c\chi_{B}I\bigr{)},$
where $\chi_{B}$ is a characteristic function on the set $B$; note the differences compared to the previous section.
Theorem 4.2.
Assume that $A_{D}$ has positive definite inclusions in $D$, corresponding to case (a) of Assumption 4.1(i). Let $A_{0}\geq\alpha I$ and assume $A_{D}\leq\beta I$ in the Loewner order in $\Omega$, for $\alpha,\beta>0$. For any measurable $B\subseteq\Omega$ we have
$B\subseteq D\quad\text{implies}\quad\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}$.
Under all of Assumption 4.1, for any open set $B\subseteq\Omega$ we have
$\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}\quad\text{implies}\quad B\subseteq D^{\bullet}.$
Proof.
The proof is given in Section 9. ∎
Theorem 4.3.
Assume that $A_{D}$ has negative definite inclusions in $D$, corresponding to case (b) of Assumption 4.1(i). For any measurable $B\subseteq\Omega$ we have
$B\subseteq D\quad\text{implies}\quad\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\leq\Lambda(A_{D})-\Lambda(A_{0}).$
Under all of Assumption 4.1, for any open set $B\subseteq\Omega$ we have
$\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\leq\Lambda(A_{D})-\Lambda(A_{0})\quad\text{implies}\quad B\subseteq D^{\bullet}.$
Proof.
The proof is given in Section 10. ∎
Remark $4.$
Under the given assumptions, for positive definite inclusions, Theorem 4.2 gives
$D^{\circ}\subseteq\cup\,\{\,B\subset\Omega\text{ open ball}\mid\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}\,\}\subset D^{\bullet},$
while for negative definite inclusions, Theorem 4.3 gives
$D^{\circ}\subseteq\cup\,\{\,B\subset\Omega\text{ open ball}\mid\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\leq\Lambda(A_{D})-\Lambda(A_{0})\,\}\subset D^{\bullet}.$
One can of course replace “open ball” by some other basis of open sets for the Euclidean topology.
5. Operator inequalities for Neumann-to-Dirichlet maps
Although we will only consider coefficients from $\mathcal{H}_{\textup{SA}}(\Omega)$ in the reconstruction methods, we still prove the following more general inequalities based on $\mathcal{H}(\Omega)$ for possible future reference. See also [24, Lemma 2.1] for the isotropic case with complex conductivity.
Theorem 5.1.
Let $A_{1},A_{2}\in\mathcal{H}(\Omega)$ and $f\in L^{2}_{\diamond}(\Gamma)$. For $j\in\{1,2\}$ we denote $\Lambda_{j}=\Lambda(A_{j})$ and $u_{j}=u_{f}^{A_{j}}$. Assume $\kappa\in\mathbb{C}$ is such that $\kappa A_{1}\in\mathcal{H}(\Omega)$. We also define
$\mathcal{B}_{\kappa}$ $=(\kappa A_{1})^{\textup{I}}[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{1})^{\textup{I}},$
$\mathcal{C}_{\kappa}$ $=(\kappa A_{2})^{\textup{I}}[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{2})^{\textup{I}}+\mathrm{i}\bigl{[}(\kappa A_{2})^{\textup{R}}[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{2})^{\textup{I}}-(\kappa A_{2})^{\textup{I}}[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{2})^{\textup{R}}\bigr{]}.$
Then
$\langle f,[\overline{\kappa}(\Lambda_{1}-\Lambda_{2})]^{\textup{R}}f\rangle$ $\geq\int_{\Omega}\bigl{[}[\kappa(A_{2}-A_{1})]^{\textup{R}}-\mathcal{B}_{\kappa}\bigr{]}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x,$ (5.1)
$\langle f,[\overline{\kappa}(\Lambda_{1}-\Lambda_{2})]^{\textup{R}}f\rangle$ $\leq\int_{\Omega}\bigl{[}(\kappa A_{2})^{\textup{R}}[(\kappa A_{1})^{\textup{R}}]^{-1}[\kappa(A_{2}-A_{1})]^{\textup{R}}+\mathcal{C}_{\kappa}\bigr{]}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.$ (5.2)
Proof.
We start by proving (5.1). A computation reveals that
$(\kappa A_{1})[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{1})^{*}=(\kappa A_{1})^{\textup{R}}+\mathcal{B}_{\kappa},$
which is used below, combined with (2.3) and (2.4):
$0$ $\leq\int_{\Omega}\bigl{\lvert}[(\kappa A_{1})^{\textup{R}}]^{1/2}\nabla u_{1}-[(\kappa A_{1})^{\textup{R}}]^{-1/2}(\kappa A_{1})^{*}\nabla u_{2}\bigr{\rvert}^{2}\,\mathrm{d}x$
$=\mathop{\textup{Re}}\Bigl{(}\kappa\int_{\Omega}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x\Bigr{)}-2\mathop{\textup{Re}}\Bigl{(}\kappa\int_{\Omega}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x\Bigr{)}$
$\phantom{={}}+\int_{\Omega}\bigl{[}(\kappa A_{1})^{\textup{R}}+\mathcal{B}_{\kappa}\bigr{]}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x$
$=\langle f,[\overline{\kappa}(\Lambda_{1}-\Lambda_{2})]^{\textup{R}}f\rangle+\int_{\Omega}\bigl{[}(\kappa A_{1})^{\textup{R}}-(\kappa A_{2})^{\textup{R}}+\mathcal{B}_{\kappa}\bigr{]}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.$
Next we prove (5.2). A computation reveals that
$(\kappa A_{2})^{*}[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{2})$ $=(\kappa A_{2})^{\textup{R}}[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{2})^{\textup{R}}+\mathcal{C}_{\kappa}$
$=(\kappa A_{2})^{\textup{R}}+(\kappa A_{2})^{\textup{R}}[(\kappa A_{1})^{\textup{R}}]^{-1}[\kappa(A_{2}-A_{1})]^{\textup{R}}+\mathcal{C}_{\kappa},$
which is used below, combined with (2.3) and (2.4):
$0$ $\leq\int_{\Omega}\bigl{\lvert}[(\kappa A_{1})^{\textup{R}}]^{1/2}\nabla u_{1}-[(\kappa A_{1})^{\textup{R}}]^{-1/2}(\kappa A_{2})\nabla u_{2}\bigr{\rvert}^{2}\,\mathrm{d}x$
$=\mathop{\textup{Re}}\Bigl{(}\kappa\int_{\Omega}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x\Bigr{)}-2\mathop{\textup{Re}}\Bigl{(}\kappa\int_{\Omega}A_{2}\nabla u_{2}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x\Bigr{)}$
$\phantom{={}}+\int_{\Omega}\bigl{[}(\kappa A_{2})^{\textup{R}}+(\kappa A_{2})^{\textup{R}}[(\kappa A_{1})^{\textup{R}}]^{-1}[\kappa(A_{2}-A_{1})]^{\textup{R}}+\mathcal{C}_{\kappa}\bigr{]}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x$
$=\langle f,[\overline{\kappa}(\Lambda_{2}-\Lambda_{1})]^{\textup{R}}f\rangle+\int_{\Omega}\bigl{[}(\kappa A_{2})^{\textup{R}}[(\kappa A_{1})^{\textup{R}}]^{-1}[\kappa(A_{2}-A_{1})]^{\textup{R}}+\mathcal{C}_{\kappa}\bigr{]}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.\qed$
Remark 5.2.
Note that in the isotropic case, the large square bracket in $\mathcal{C}_{\kappa}$ vanishes (after the imaginary unit). Also, if $\kappa A_{1}\in\mathcal{H}_{\textup{SA}}(\Omega)$ then $\mathcal{B}_{\kappa}=0$ and if $\kappa A_{2}\in\mathcal{H}_{\textup{SA}}(\Omega)$ then $\mathcal{C}_{\kappa}=0$.
We get some corollaries of Theorem 5.1, based on the choice of $\kappa$.
Corollary 5.3.
In the setting of Theorem 5.1 with $\kappa=1$, we have
$\langle f,(\Lambda^{\!\textup{R}}_{1}-\Lambda^{\!\textup{R}}_{2})f\rangle$ $\geq\int_{\Omega}\bigl{[}A^{\textup{R}}_{2}-A^{\textup{R}}_{1}-\mathcal{B}_{1}\bigr{]}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x,$
$\langle f,(\Lambda^{\!\textup{R}}_{1}-\Lambda^{\!\textup{R}}_{2})f\rangle$ $\leq\int_{\Omega}\bigl{[}A^{\textup{R}}_{2}(A^{\textup{R}}_{1})^{-1}(A^{\textup{R}}_{2}-A^{\textup{R}}_{1})+\mathcal{C}_{1}\bigr{]}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.$
Corollary 5.4.
In the setting of Theorem 5.1 with $\kappa=-\mathrm{i}$, provided that $A^{\textup{I}}_{1}\in\mathcal{H}_{\textup{SA}}(\Omega)$, we have
$\langle f,(\Lambda^{\!\textup{I}}_{2}-\Lambda^{\!\textup{I}}_{1})f\rangle$ $\geq\int_{\Omega}\bigl{[}A^{\textup{I}}_{2}-A^{\textup{I}}_{1}-\mathcal{B}_{-\mathrm{i}}\bigr{]}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x,$
$\langle f,(\Lambda^{\!\textup{I}}_{2}-\Lambda^{\!\textup{I}}_{1})f\rangle$ $\leq\int_{\Omega}\bigl{[}A^{\textup{I}}_{2}(A^{\textup{I}}_{1})^{-1}(A^{\textup{I}}_{2}-A^{\textup{I}}_{1})+\mathcal{C}_{-\mathrm{i}}\bigr{]}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.$
What we will need for the monotonicity method is the following, corresponding to the special case of Theorem 5.1 where $A_{1},A_{2}\in\mathcal{H}_{\textup{SA}}(\Omega)$ and $\kappa=1$. We also provide a slightly different proof that is arguably easier to read.
Proposition 5.5.
Let $A_{1},A_{2}\in\mathcal{H}_{\textup{SA}}(\Omega)$ and $f\in L^{2}_{\diamond}(\Gamma)$. For $j\in\{1,2\}$ we denote $\Lambda_{j}=\Lambda(A_{j})$ and $u_{j}=u_{f}^{A_{j}}$. Then
$\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle$ $\geq\int_{\Omega}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x,$ (5.3)
$\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle$ $\leq\int_{\Omega}A_{2}A_{1}^{-1}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.$ (5.4)
Proof.
We start by proving (5.3), which is by the computation below combined with (2.3),
$\int_{\Omega}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x$
$\hskip 42.67912pt\leq\int_{\Omega}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x+\int_{\Omega}\lvert A_{1}^{1/2}\nabla(u_{1}-u_{2})\rvert^{2}\,\mathrm{d}x$
$\hskip 42.67912pt=\int_{\Omega}A_{2}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x+\int_{\Omega}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x-2\mathop{\textup{Re}}\Bigl{(}\int_{\Omega}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x\Bigr{)}$
$\hskip 42.67912pt=\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle.$
Next we prove (5.4). Using that
$A_{2}A_{1}^{-1}(A_{2}-A_{1})=A_{2}A_{1}^{-1}A_{2}-A_{2}$
is self-adjoint, we get
$\int_{\Omega}A_{2}A_{1}^{-1}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x$
$\hskip 42.67912pt\geq\int_{\Omega}A_{2}A_{1}^{-1}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x-\int_{\Omega}\lvert A_{1}^{1/2}\nabla u_{1}-A_{1}^{-1/2}A_{2}\nabla u_{2}\rvert^{2}\,\mathrm{d}x$
$\hskip 42.67912pt=-\int_{\Omega}A_{2}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x-\int_{\Omega}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x+2\mathop{\textup{Re}}\Bigl{(}\int_{\Omega}A_{2}\nabla u_{2}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x\Bigr{)}$
$\hskip 42.67912pt=\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle.\qed$
Note that the first inequality in Proposition 5.5 implies monotonicity of the ND maps, since
$\int_{\Omega}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x\leq\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle\leq\int_{\Omega}(A_{2}-A_{1})\nabla u_{1}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x.$ (5.5)
It is not immediately clear how one would make use of the second inequality in Proposition 5.5, due to the matrix product in the front. What we need to consider is when $A_{2}-A_{1}$ is positive/negative (semi)definite, if we can also infer that $A_{2}A_{1}^{-1}(A_{2}-A_{1})$ has the same properties. This is the result of the following proposition.
Proposition 5.6.
Let $c\geq 0$ and $V\subseteq\Omega$. Let $A_{1},A_{2}\in\mathcal{H}_{\textup{SA}}(\Omega)$ with $A_{1}\leq\beta I$ and $A_{2}\geq\alpha I$ in $V$ for $\alpha,\beta>0$.
  1. (i)
    If $A_{2}-A_{1}\geq cI$ in $V$, then $A_{2}A_{1}^{-1}(A_{2}-A_{1})\geq cI$ in $V$.
  2. (ii)
    If $A_{2}-A_{1}\leq-cI$ in $V$, then $A_{2}A_{1}^{-1}(A_{2}-A_{1})\leq-c(\tfrac{\alpha}{\beta})^{2}I$ in $V$.
Proof.
Let $A_{2}-A_{1}\geq cI$ in the Loewner order in $V$ for some $c\geq 0$. Let $\xi\in\mathbb{C}^{d}$ and note that $A_{1}^{-1}$ is positive definite. Now we put the pieces together:
$A_{2}A_{1}^{-1}(A_{2}-A_{1})\xi\cdot\overline{\xi}$ $=(A_{2}-A_{1})A_{1}^{-1}(A_{2}-A_{1})\xi\cdot\overline{\xi}+(A_{2}-A_{1})\xi\cdot\overline{\xi}$
$=A_{1}^{-1}(A_{2}-A_{1})\xi\cdot\overline{(A_{2}-A_{1})\xi}+(A_{2}-A_{1})\xi\cdot\overline{\xi}$
$\geq c\lvert\xi\rvert^{2},$
which concludes the proof of (i).
For (ii), we first prove the assertion that if
$A_{1}-A_{2}\geq cI$ (5.6)
for some $c\geq 0$, then
$A_{2}^{-1}-A_{1}^{-1}\geq c\beta^{-2}I.$ (5.7)
The semidefinite case ($c=0$) of (5.7) is a simple consequence of $t\mapsto-t^{-1}$ being an operator monotone function, c.f. [8, chapter 7]. To prove (5.7), first note that we have
$A_{2}^{-1}-A_{1}^{-1}$ $=A_{1}^{-1}(A_{1}-A_{2})A_{2}^{-1}$
$=A_{1}^{-1}(A_{1}-A_{2})A_{1}^{-1}+A_{1}^{-1}(A_{1}-A_{2})(A_{2}^{-1}-A_{1}^{-1})$
$=A_{1}^{-1}\bigl{[}A_{1}-A_{2}+(A_{1}-A_{2})A_{2}^{-1}(A_{1}-A_{2})\bigr{]}A_{1}^{-1}.$
Note that $A_{1}^{-1}\geq\beta^{-1}I$ and likewise that $A_{2}^{-1}$ is positive definite. Hence, we have
$(A_{2}^{-1}-A_{1}^{-1})\xi\cdot\overline{\xi}$ $=(A_{1}-A_{2})A_{1}^{-1}\xi\cdot\overline{A_{1}^{-1}\xi}+A_{2}^{-1}(A_{1}-A_{2})A_{1}^{-1}\xi\cdot\overline{(A_{1}-A_{2})A_{1}^{-1}\xi}$
$\geq c\lvert A_{1}^{-1}\xi\rvert^{2}\geq c\beta^{-2}\lvert\xi\rvert^{2}.$
Now assume that in $V$ we have (5.6) for some $c\geq 0$. Then (5.7) implies
$A_{2}A_{1}^{-1}(A_{1}-A_{2})\xi\cdot\overline{\xi}=(A_{2}^{-1}-A_{1}^{-1})A_{2}\xi\cdot\overline{A_{2}\xi}\geq c\beta^{-2}\lvert A_{2}\xi\rvert^{2}\geq c(\tfrac{\alpha}{\beta})^{2}\lvert\xi\rvert^{2},$
which concludes the proof of (ii). ∎
6. Localized potentials for anisotropic coefficients
We extend localization results from [20] to anisotropic coefficients in $\mathcal{H}(\Omega)$. For a measurable set $V\subseteq\overline{\Omega}$ and $F\in L^{2}(V)^{d}$, we define $w=w_{F}^{A}$ as the unique solution in $H^{1}_{\diamond}(\Omega)$ to the variational problem
$\int_{\Omega}A^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x=\int_{V}F\cdot\overline{\nabla v}$ (6.1)
for all $v\in H^{1}_{\diamond}(\Omega).$ We define the virtual measurement operator $L_{V}(A)\colon L^{2}(V)^{d}\to L^{2}_{\diamond}(\Gamma)$ as
$L_{V}(A)F=w_{F}^{A}|_{\Gamma}.$
In the following, $R$ is used to denote the range of an operator.
Proposition 6.1.
For $A\in\mathcal{H}(\Omega)$ and measurable sets $V,V_{1},V_{2}\subseteq\overline{\Omega}$, there are the following properties of the virtual measurement operator:
  1. (i)
    $R(L_{V}(A))$ is independent of the function values $A|_{V}$,
  2. (ii)
    $V_{1}\subseteq V_{2}$ implies $R(L_{V_{1}}(A))\subseteq R(L_{V_{2}}(A))$,
  3. (iii)
    The adjoint $L_{V}(A)^{*}\colon L^{2}_{\diamond}(\Gamma)\to L^{2}(V)^{d}$ is given by $L_{V}(A)^{*}f=\nabla u_{f}^{A}|_{V}$.
Proof.
(i): Let $A_{1},A_{2}\in\mathcal{H}(\Omega)$ such that $A_{1}=A_{2}$ outside $V$. We will prove that $R(L_{V}(A_{1}))=R(L_{V}(A_{2}))$. Let $\varphi\in R(L_{V}(A_{1}))$, i.e. $\varphi=w|_{\Gamma}$ with $w=w_{F_{1}}^{A_{1}}$ for some $F_{1}\in L^{2}(V)^{d}$. Now define $F_{2}=(A_{2}-A_{1})^{*}\nabla w|_{V}\in L^{2}(V)^{d}$. Then (6.1) gives
$\int_{\Omega}A_{2}^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x$ $=\int_{\Omega}(A_{2}-A_{1})^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x+\int_{\Omega}A_{1}^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x$
$=\int_{V}(A_{2}-A_{1})^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x+\int_{V}F_{1}\cdot\overline{\nabla v}\,\mathrm{d}x$
$=\int_{V}(F_{1}+F_{2})\cdot\overline{\nabla v}\,\mathrm{d}x$
for all $v\in H^{1}_{\diamond}(\Omega)$. Hence, we have that $\varphi=L_{V}(A_{2})(F_{1}+F_{2})$ so indeed $\varphi\in R(L_{V}(A_{2}))$. Repeating the above proof with $A_{1}$ and $A_{2}$ interchanged results in $R(L_{V}(A_{1}))=R(L_{V}(A_{2}))$ as desired.
(ii): Assume $V_{1}\subseteq V_{2}$, and let $\varphi=L_{V_{1}}(A)F=w|_{\Gamma}$ for some $F\in L^{2}(V_{1})^{d}$. Let $\widetilde{F}$ denote the extension by zero of $F$ on the set $V_{2}$. Then
$\int_{\Omega}A^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x=\int_{V_{1}}F\cdot\overline{\nabla v}\,\mathrm{d}x=\int_{V_{2}}\widetilde{F}\cdot\overline{\nabla v}\,\mathrm{d}x$
for all $v\in H^{1}_{\diamond}(\Omega)$. Hence, we also have $\varphi=L_{V_{2}}(A)\widetilde{F}$, i.e. $R(L_{V_{1}}(A))\subseteq R(L_{V_{2}}(A))$.
(iii): Let $F\in L^{2}(V)^{d}$ and $f\in L^{2}_{\diamond}(\Gamma)$, then from the variational forms (6.1) and (2.2), for $w=w_{F}^{A}$ and $u=u_{f}^{A}$, we obtain
$\langle L_{V}(A)^{*}f,F\rangle_{L^{2}(V)^{d}}$ $=\langle f,L_{V}(A)F\rangle_{L^{2}(\Gamma)}=\langle f,w\rangle_{L^{2}(\Gamma)}$
$=\int_{\Omega}A\nabla u\cdot\overline{\nabla w}\,\mathrm{d}x=\int_{V}\overline{F}\cdot\nabla u\,\mathrm{d}x=\langle\nabla u|_{V},F\rangle_{L^{2}(V)^{d}}.$
Since this holds for all $F\in L^{2}(V)^{d}$, we conclude that $L_{V}(A)^{*}f=\nabla u|_{V}$. ∎
We note that range inclusions imply norm inequalities for the adjoint operators.
Lemma 6.2 (Lemma 2.5 in [20]).
Let $H$, $K_{1}$, and $K_{2}$ be Hilbert spaces and let $T_{j}\in\mathscr{L}(K_{j},H)$ for $j\in\{1,2\}$. Then
$R(T_{1})\subseteq R(T_{2})\qquad\text{if and only if}\qquad\exists C>0,\,\forall x\in H\colon\lVert T_{1}^{*}x\rVert_{K_{1}}\leq C\lVert T_{2}^{*}x\rVert_{K_{2}}.$
Finally, we are able to prove the existence of localized potentials.
Theorem 6.3.
Let $U\subset\overline{\Omega}$ be a relatively open connected set that intersects $\Gamma$ and let $B\subset U$ be a non-empty open set. Assume $A\in\mathcal{H}(\Omega)$ where both $A$ and $A^{*}$ satisfy the UCP in $U$. Then there are sequences $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ and $(u_{j})$ in $H^{1}_{\diamond}(\Omega)$, with $u_{j}=u_{f_{j}}^{A}$, such that
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
Proof.
We let $V_{1}\subseteq B\subset V\subseteq U$ where $V_{1}$ and $V$ are relatively open non-empty connected sets, and define $V_{2}=\Omega\setminus\overline{V}$. These sets are chosen such that $V$ intersects $\Gamma$, $\mathop{\textup{dist}}(V_{1},V_{2})>0$, and that $\partial(\Omega\setminus\overline{V_{j}})$ is Lipschitz continuous.
We start by proving that $R(L_{V_{1}}(A))$ and $R(L_{V_{2}}(A))$ have a trivial intersection. Hence, let $\varphi\in R(L_{V_{1}}(A))\cap R(L_{V_{2}}(A))$, i.e. $\varphi=w_{j}|_{\Gamma}$ with $w_{j}=w_{F_{j}}^{A}$ for some $F_{j}\in L^{2}(V_{j})^{d}$ for $j\in\{1,2\}$. From (6.1) we conclude that
$\int_{\Omega}A^{*}\nabla w_{j}\cdot\overline{\nabla v}\,\mathrm{d}x=0\text{ for all }v\in H^{1}(\Omega\setminus\overline{V_{j}}).$
From the weak form of the conductivity problem (2.2), with conductivity $A^{*}$, we thus have
$-\nabla\cdot(A^{*}\nabla w_{j})$ $=0\text{ in }\Omega\setminus\overline{V_{j}}$
$\nu\cdot(A^{*}\nabla w_{j})$ $=0\text{ on }\partial(\Omega\setminus\overline{V_{j}}),$
in particular $\nu\cdot(A^{*}\nabla w_{j})\equiv 0$ on $\partial V\cap\Gamma$, i.e. the Cauchy data of $w_{1}$ and $w_{2}$ coincide on $\partial V\cap\Gamma$. Unique continuation entails that $w_{1}=w_{2}$ in $V\setminus\overline{V_{1}}$. We may now “glue” $w_{1}$ and $w_{2}$ together, using that they agree in $V\setminus\overline{V_{1}}$:
$u=\begin{dcases}w_{1}&\text{in }\Omega\setminus\overline{V_{1}},\\ w_{2}&\text{in }V_{1}.\end{dcases}$
Now we have $\nabla\cdot(A^{*}\nabla u)=0$ in $\Omega$ and $\nu\cdot(A^{*}\nabla u)=\nu\cdot(A^{*}\nabla w_{1})\equiv 0$ on $\partial\Omega$. By uniqueness of such a solution in $H^{1}_{\diamond}(\Omega)$, it follows that $u\equiv 0$ in $\Omega$, so in particular $\varphi=w_{1}|_{\Gamma}=u|_{\Gamma}\equiv 0$, i.e.
$R(L_{V_{1}}(A))\cap R(L_{V_{2}}(A))=\{0\}.$
Next we prove that $R(L_{V_{1}}(A))$ is non-trivial. Suppose that $0\equiv L_{V_{1}}(A)^{*}f=\nabla u_{f}^{A}|_{V_{1}}$, then $u_{f}^{A}$ is a constant function in $V_{1}$, and also in $V$ by unique continuation. In particular, $f=\nu\cdot(A\nabla u_{f}^{A})|_{\Gamma}\equiv 0$ on $\partial V\cap\Gamma$. Thus, $L_{V_{1}}(A)^{*}$ is injective on
$W=\{\,f\in L^{2}_{\diamond}(\Gamma)\mid\mathop{\textup{supp}}f\subseteq\partial V\cap\Gamma\,\}.$
Since $\overline{R(L_{V_{1}}(A))}=N(L_{V_{1}}(A)^{*})^{\perp}$ then $R(L_{V_{1}}(A))$ contains a dense subset of $W$ and must be non-trivial.
In total we have that $R(L_{V_{1}}(A))\not\subseteq R(L_{V_{2}}(A))$. From Lemma 6.2 this implies the existence of a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A}$ satisfy (c.f. Proposition 6.1(iii)):
$\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x\geq\int_{V_{1}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\lVert L_{V_{1}}(A)^{*}f_{j}\rVert_{L^{2}(V_{1})^{d}}^{2}\to\infty$
and
$\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x\leq\int_{V_{2}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\lVert L_{V_{2}}(A)^{*}f_{j}\rVert_{L^{2}(V_{2})^{d}}^{2}\to 0$
for $j\to\infty$. ∎
The following result shows that the sequence $(f_{j})$ of Neumann conditions (e.g. from Theorem 6.3) such that the corresponding electric power $\lvert\nabla u_{j}\rvert^{2}$ is localized, can be chosen independently of the function values of $A$ on the set where the power tends to zero.
Proposition 6.4.
Let $B$ and $U$ be as in Theorem 6.3, and assume for $A_{1}\in\mathcal{H}(\Omega)$ and $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ that $u_{j}=u_{f_{j}}^{A_{1}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
If $A_{2}\in\mathcal{H}(\Omega)$ with $\mathop{\textup{supp}}(A_{1}-A_{2})\subseteq\overline{\Omega}\setminus U$, then $\widehat{u}_{j}=u_{f_{j}}^{A_{2}}$ also satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x=0.$
Proof.
Let $V_{1}=B$ and $V_{2}=\overline{\Omega}\setminus U$. From Proposition 6.1(iii) it amounts to proving
$\lim_{j\to\infty}\lVert L_{V_{1}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{1})^{d}}=\infty\quad\text{and}\quad\lim_{j\to\infty}\lVert L_{V_{2}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{2})^{d}}=0.$´
From Proposition 6.1(i) we have that $R(L_{V_{2}}(A_{1}))=R(L_{V_{2}}(A_{2}))$. Hence a combination of Proposition 6.1(iii) and Lemma 6.2 gives
$\lVert L_{V_{2}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{2})^{d}}\leq C\lVert L_{V_{2}}(A_{1})^{*}f_{j}\rVert_{L^{2}(V_{2})^{d}}\to 0\text{ for }j\to\infty.$
By the same argumentation, we also have $R(L_{V_{1}\cup V_{2}}(A_{1}))=R(L_{V_{1}\cup V_{2}}(A_{2}))$, which implies
$\lVert L_{V_{1}\cup V_{2}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{1}\cup V_{2})^{d}}$ $\geq c\lVert L_{V_{1}\cup V_{2}}(A_{1})^{*}f_{j}\rVert_{L^{2}(V_{1}\cup V_{2})^{d}}$
$\geq c\lVert L_{V_{1}}(A_{1})^{*}f_{j}\rVert_{L^{2}(V_{1})^{d}}\to\infty\text{ for }j\to\infty.$
Finally we combine the two above limits, using Proposition 6.1(iii),
$\lVert L_{V_{1}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{1})^{d}}^{2}=\lVert L_{V_{1}\cup V_{2}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{1}\cup V_{2})^{d}}^{2}-\lVert L_{V_{2}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{2})^{d}}^{2}\to\infty\text{ for }j\to\infty,$
thereby concluding the proof. ∎
7. Proof of Theorem 3.6
The proof that $D\subseteq C$ implies $\Lambda^{-}_{C}\geq\Lambda(A_{D})\geq\Lambda^{+}_{C}$ follows directly from Assumption 3.4(i) and (5.5), since we have
$A^{-}_{C}\leq A_{D}\leq A^{+}_{C}$
in the Loewner order in $\Omega$.
Now assume that $D\not\subseteq C$, i.e. that $D^{\bullet}\not\subseteq C$. Since both sets are closures of opens sets and have connected complements, there is a relatively open set that connects $D^{\bullet}\setminus C$ with $\Gamma$ that we may use for localization. Specifically, there exist a relatively open connected set $U\subset\overline{\Omega}\setminus C$ that intersects $\Gamma$ and an open ball $B\subset D^{\bullet}\cap U$. Moreover, Assumption 3.4(iii) ensures that we can pick $B$ and $U$ such that we arrive at one of two cases:
  • Case 1: $A_{D}-A_{0}$ is positive semidefinite in $U$ and uniformly positive definite in $B$.
  • Case 2: $A_{D}-A_{0}$ is negative semidefinite in $U$ and uniformly negative definite in $B$.
In the first case we will prove that $\Lambda(A_{D})\not\geq\Lambda^{+}_{C}$, and in the second case we will prove that $\Lambda^{-}_{C}\not\geq\Lambda(A_{D})$. Together the two cases prove the contrapositive formulation of the assertion
$\Lambda^{-}_{C}\geq\Lambda(A_{D})\geq\Lambda^{+}_{C}\quad\text{implies}\quad D\subseteq C.$
Case 1: Using Theorem 6.3 and Proposition 6.4 (as $C\subset\Omega\setminus U$), there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A^{+}_{C}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
Note that $A_{C}^{+}=A_{0}$ in $U$ where $A_{D}-A_{0}$ is positive semidefinite, and also in $B$ where $A_{D}-A_{0}$ is uniformly positive definite. From Proposition 5.6(ii) then $A_{0}A_{D}^{-1}(A_{0}-A_{D})$ is negative semidefinite in $U$, and there exists a scalar $\tilde{c}>0$ such that
$A_{0}A_{D}^{-1}(A_{0}-A_{D})\leq-\tilde{c}I\text{ in }B.$
Now Proposition 5.5 gives
$\langle f_{j},\bigl{[}\Lambda(A_{D})-\Lambda^{+}_{C}\bigr{]}f_{j}\rangle$ $\leq\int_{\Omega}A^{+}_{C}A_{D}^{-1}(A^{+}_{C}-A_{D})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x$
$\leq\lVert A^{+}_{C}A_{D}^{-1}(A^{+}_{C}-A_{D})\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x-\tilde{c}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x.$
We conclude that
$\lim_{j\to\infty}\langle f_{j},\bigl{[}\Lambda(A_{D})-\Lambda^{+}_{C}\bigr{]}f_{j}\rangle=-\infty,$
in particular, $\Lambda(A_{D})\not\geq\Lambda^{+}_{C}$.
Case 2: Using Theorem 6.3 and Proposition 6.4 (as $C\subset\Omega\setminus U$), there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A^{-}_{C}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
Note that $A_{C}^{-}=A_{0}$ in $U$ where $A_{D}-A_{0}$ is negative semidefinite, and also in $B$ where $A_{D}-A_{0}$ is uniformly negative definite. From Proposition 5.5 we have
$\langle f_{j},\bigl{[}\Lambda^{-}_{C}-\Lambda(A_{D})\bigr{]}f_{j}\rangle$ $\leq\int_{\Omega}(A_{D}-A^{-}_{C})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x$
$\leq\lVert A_{D}-A^{-}_{C}\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x-c\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x$
for a scalar $c>0$. We conclude that
$\lim_{j\to\infty}\langle f_{j},\bigl{[}\Lambda^{-}_{C}-\Lambda(A_{D})\bigr{]}f_{j}\rangle=-\infty,$
in particular, $\Lambda^{-}_{C}\not\geq\Lambda(A_{D})$.
8. Proof of Theorem 3.7
Assume $D\subseteq C$ and let $u_{0}=u_{f}^{A_{0}}$. From Proposition 5.5 we have
$\langle f,\bigl{[}\Lambda(A_{D})-\Lambda(A_{0})-\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\bigr{]}f\rangle$ $\geq\int_{\Omega}(A_{0}-A_{D})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x+\int_{C}(\beta I-A_{0})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x$
$=\int_{D}(A_{0}-A_{D})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x+\int_{C}(\beta I-A_{0})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x$
$\geq\int_{C\setminus D}(\beta I-A_{0})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x\geq 0.$
For the next part we need an additional bound. Let $\xi\in\mathbb{C}^{d}$, then we have
$A_{0}A_{D}^{-1}A_{0}\xi\cdot\overline{\xi}=A_{D}^{-1}A_{0}\xi\cdot\overline{A_{0}\xi}\leq\alpha^{-1}\lvert A_{0}\xi\rvert^{2}\leq\tfrac{\beta^{2}}{\alpha}\lvert\xi\rvert^{2}.$ (8.1)
Note that $A_{0}\leq\beta I\leq\frac{\beta^{2}}{\alpha}I$. Now from Proposition 5.5 and (8.1) we get
$\langle f,\bigl{[}\Lambda(A_{0})-\Lambda(A_{D})+\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\bigr{]}f\rangle$
$\hskip-85.35826pt\geq\int_{\Omega}A_{0}A_{D}^{-1}(A_{D}-A_{0})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x+\int_{C}\bigl{(}\tfrac{\beta^{2}}{\alpha}I-A_{0}\bigr{)}\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x$
$\hskip-85.35826pt=\int_{D}(A_{0}-A_{0}A_{D}^{-1}A_{0})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x+\int_{C}\bigl{(}\tfrac{\beta^{2}}{\alpha}I-A_{0}\bigr{)}\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x$
$\hskip-85.35826pt\geq\int_{C\setminus D}\bigl{(}\tfrac{\beta^{2}}{\alpha}I-A_{0}\bigr{)}\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x\geq 0.$
This concludes the proof for the first part of the theorem statement.
Now assume that $D\not\subseteq C$, i.e. that $D^{\bullet}\not\subseteq C$. Since both sets are closures of opens sets and have connected complements, there is a relatively open set that connects $D^{\bullet}\setminus C$ with $\Gamma$ that we may use for localization. Specifically, there exist a relatively open connected set $U\subset\overline{\Omega}\setminus C$ that intersects $\Gamma$ and an open ball $B\subset D^{\bullet}\cap U$. Moreover, Assumption 3.4(iii) ensures that we can pick $B$ and $U$ such that we arrive at one of two cases:
  • Case 1: $A_{D}-A_{0}$ is positive semidefinite in $U$ and uniformly positive definite in $B$.
  • Case 2: $A_{D}-A_{0}$ is negative semidefinite in $U$ and uniformly negative definite in $B$.
In the first case we will prove that $\Lambda(A_{D})-\Lambda(A_{0})\not\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}$, and in the second case we will prove that $\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\not\geq\Lambda(A_{D})-\Lambda(A_{0})$. Together the two cases prove the contrapositive formulation of the assertion
$\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\geq\Lambda(A_{D})-\Lambda(A_{0})\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\quad\text{implies}\quad D\subseteq C.$
For both cases we use the same sequence of localized potentials. From Theorem 6.3 there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A_{0}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
Case 1: We now combine Proposition 5.5 with Proposition 5.6(ii), using the facts that $A_{D}-A_{0}$ is positive semidefinite in $U$ and $A_{D}-A_{0}\geq cI$ in $B$ for some scalar $c>0$, which results in
$\langle f_{j},\bigl{[}\Lambda(A_{D})-\Lambda(A_{0})-\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\bigr{]}f_{j}\rangle$
$\hskip-113.81102pt\leq\int_{\Omega}A_{0}A_{D}^{-1}(A_{0}-A_{D})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x+\int_{C}(\beta I-A_{0})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x$
$\hskip-113.81102pt\leq\lVert A_{0}A_{D}^{-1}(A_{0}-A_{D})\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x-c\bigl{(}\tfrac{\alpha}{\beta}\bigr{)}^{2}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x+\lVert\beta I-A_{0}\rVert_{*}\int_{C}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x.$
Since $C\subset\Omega\setminus U$, we conclude that
$\lim_{j\to\infty}\langle f_{j},\bigl{[}\Lambda(A_{D})-\Lambda(A_{0})-\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\bigr{]}f_{j}\rangle=-\infty,$
in particular, $\Lambda(A_{D})-\Lambda(A_{0})\not\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}$.
Case 2: We use Proposition 5.5 together with the facts that $A_{D}-A_{0}$ is negative semidefinite in $U$ and $A_{D}-A_{0}\leq-cI$ in $B$ for some scalar $c>0$, which results in
$\langle f_{j},\bigl{[}\Lambda(A_{0})-\Lambda(A_{D})+\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\bigr{]}f_{j}\rangle$
$\hskip-113.81102pt\leq\int_{\Omega}(A_{D}-A_{0})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x+\int_{C}\bigl{(}\tfrac{\beta^{2}}{\alpha}I-A_{0}\bigr{)}\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x$
$\hskip-113.81102pt\leq\lVert A_{D}-A_{0}\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x-c\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x+\lVert\tfrac{\beta^{2}}{\alpha}I-A_{0}\rVert_{*}\int_{C}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x.$
Since $C\subset\Omega\setminus U$, we conclude that
$\lim_{j\to\infty}\langle f_{j},\bigl{[}\Lambda(A_{0})-\Lambda(A_{D})+\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\bigr{]}f_{j}\rangle=-\infty,$
in particular, $\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\not\geq\Lambda(A_{D})-\Lambda(A_{0})$.
9. Proof of Theorem 4.2
Assume $B\subseteq D$ and let $u_{0}=u_{f}^{A_{0}}$. By assumption, $A_{D}-A_{0}\geq cI$ in $D$ for some $c>0$, from which Proposition 5.6(ii) implies that
$A_{0}A_{D}^{-1}(A_{D}-A_{0})\geq c\bigl{(}\tfrac{\alpha}{\beta}\bigr{)}^{2}I\text{ in }D.$
By Proposition 5.5, we have
$\langle f,\bigl{[}\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}+\Lambda(A_{0})-\Lambda(A_{D})\bigr{]}f\rangle$ $\geq-c\bigl{(}\tfrac{\alpha}{\beta}\bigr{)}^{2}\int_{B}\lvert\nabla u_{0}\rvert^{2}\,\mathrm{d}x+\int_{\Omega}A_{0}A_{D}^{-1}(A_{D}-A_{0})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x$
$\geq c\bigl{(}\tfrac{\alpha}{\beta}\bigr{)}^{2}\int_{D\setminus B}\lvert\nabla u_{0}\rvert^{2}\,\mathrm{d}x\geq 0.$
Next, assume $B\not\subseteq D^{\bullet}$. Since $D^{\bullet}$ is closed and has connected complement, while $B$ is a non-empty open set, there exist a relatively open and connected set $U\subseteq\overline{\Omega}\setminus D^{\bullet}$ intersecting $\Gamma$ and $B$, and an open ball $\widehat{B}\subset U\cap B$. From Theorem 6.3 there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A_{0}}$ satisfy
$\lim_{j\to\infty}\int_{\widehat{B}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
Since $D\subset\Omega\setminus U$, Proposition 5.5 gives
$\langle f_{j},\bigl{[}\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}+\Lambda(A_{0})-\Lambda(A_{D})\bigr{]}f_{j}\rangle$ $\leq-c\bigl{(}\tfrac{\alpha}{\beta}\bigr{)}^{2}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x+\int_{\Omega}(A_{D}-A_{0})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x$
$\leq-c\bigl{(}\tfrac{\alpha}{\beta}\bigr{)}^{2}\int_{\widehat{B}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x+\lVert A_{D}-A_{0}\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x.$
We conclude that
$\lim_{j\to\infty}\langle f_{j},\bigl{[}\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}+\Lambda(A_{0})-\Lambda(A_{D})\bigr{]}f_{j}\rangle=-\infty,$
in particular, $\Lambda(A_{D})-\Lambda(A_{0})\not\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}$. This proves the contrapositive formulation of the assertion
$\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}\quad\text{implies}\quad B\subseteq D^{\bullet}.$
10. Proof of Theorem 4.3
Assume $B\subseteq D$ and let $u_{0}=u_{f}^{A_{0}}$. By assumption, $A_{0}-A_{D}\geq cI$ in $D$, so Proposition 5.5 yields
$\langle f,\bigl{[}\Lambda(A_{D})-\Lambda(A_{0})-\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\bigr{]}f\rangle$ $\geq-c\int_{B}\lvert\nabla u_{0}\rvert^{2}\,\mathrm{d}x+\int_{\Omega}(A_{0}-A_{D})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x$
$\geq c\int_{D\setminus B}\lvert\nabla u_{0}\rvert^{2}\,\mathrm{d}x\geq 0.$
Next, assume $B\not\subseteq D^{\bullet}$. Since $D^{\bullet}$ is closed and has connected complement, while $B$ is a non-empty open set, there exist a relatively open and connected set $U\subseteq\overline{\Omega}\setminus D^{\bullet}$ intersecting $\Gamma$ and $B$, and an open ball $\widehat{B}\subset U\cap B$. From Theorem 6.3 there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A_{0}}$ satisfy
$\lim_{j\to\infty}\int_{\widehat{B}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
Since $D\subset\Omega\setminus U$, Proposition 5.5 gives
$\langle f_{j},\bigl{[}\Lambda(A_{D})-\Lambda(A_{0})-\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\bigr{]}f_{j}\rangle$ $\leq-c\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x+\int_{\Omega}A_{0}A_{D}^{-1}(A_{0}-A_{D})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x$
$\leq-c\int_{\widehat{B}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x+\lVert A_{0}A_{D}^{-1}(A_{0}-A_{D})\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x.$
We conclude that
$\lim_{j\to\infty}\langle f_{j},\bigl{[}\Lambda(A_{D})-\Lambda(A_{0})-\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\bigr{]}f_{j}\rangle=-\infty,$
in particular, $\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\not\leq\Lambda(A_{D})-\Lambda(A_{0})$. This proves the contrapositive formulation of the assertion
$\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\leq\Lambda(A_{D})-\Lambda(A_{0})\quad\text{implies}\quad B\subseteq D^{\bullet}.$
Acknowledgements
The authors are supported by grant 10.46540/3120-00003B from Independent Research Fund Denmark.