### v2

# Direct reconstruction of anisotropic self-adjoint inclusions in the Calderón problem

Abstract.
We extend the monotonicity method for direct exact reconstruction of inclusions in the partial data Calderón problem, to the case of general anisotropic conductivities in any spatial dimension $d\geq 2$. From a local Neumann-to-Dirichlet map, we give reconstruction methods of inclusions based on unknown anisotropic self-adjoint perturbations to a known anisotropic conductivity coefficient. This additionally provides new insights into the non-uniqueness issues of the anisotropic Calderón problem.
The main assumption is a definiteness condition for the perturbations near the outer inclusion boundaries. Beyond this condition, they are $L^{\infty}$-perturbations that may be indefinite away from the outer inclusion boundaries, and with no boundary regularity requirement for the inclusions. Alternatively, we allow extreme parts that are perfectly insulating or perfectly conducting, in which case we require Lipschitz regularity of the outer inclusion boundaries.
Keywords: anisotropic Calderón problem, inclusion detection, monotonicity method.
2020 Mathematics Subject Classification: 35R30, 35R05, 47H05.
1. Introduction
The anisotropic Calderón problem, the inverse problem of determining an elliptic matrix-valued conductivity coefficient $A$ from boundary measurements of the conductivity equation
[…]
is of major interest in the inverse problems community. Unlike the isotropic (scalar-valued) case, the anisotropic Calderón problem suffers from significant non-uniqueness issues. Indeed if boundary measurements are taken on $\Gamma\subseteq\partial\Omega$, for any $H^{1}$-diffeomorphism $\Phi\colon\overline{\Omega}\to\overline{\Omega}$ for which $\Phi|_{\Gamma}$ is the identity, the conductivity coefficient defined via the push-forward
$\widetilde{A}=\frac{\mathop{\textup{D}\Phi}\nolimits A\mathop{\textup{D}\Phi}\nolimits^{\textup{T}}}{\lvert\det\mathop{\textup{D}\Phi}\nolimits\rvert}\circ\Phi^{-1}$ (1.1)
produces the same local Neumann-to-Dirichlet (ND) map as the coefficient $A$. For $\Gamma=\partial\Omega$ and with spatial dimension $d=2$, the obstructions to uniqueness are characterized by (1.1) for real-valued symmetric $L^{\infty}$-conductivities [4, 38]. In spatial dimension $d\geq 3$, the full extent of the non-uniqueness issues (beyond real-analytic coefficients) remains one of the most important unanswered questions in the inverse problems community [33, 40]. In a fractional Calderón problem, using the stronger non-local properties, the obstructions to uniqueness are characterized for $d\geq 2$ [10].
Despite knowledge on the possible obstructions, the non-uniqueness has prevented development of practically useful reconstruction methods. Unique recovery is only guaranteed with very strong restrictions on the considered class of coefficients [2, 3, 11].
In this paper we study the exact reconstruction of anisotropic inclusions in any spatial dimension $d\geq 2$. We reconstruct the outer shape $D^{\bullet}$ of
$D=\mathop{\textup{supp}}(A_{D}-A_{0}),$
from knowledge of $A_{0}$ and the local ND map $\Lambda(A_{D})$ associated with the unknown conductivity $A_{D}$. Here $D^{\bullet}$ is the smallest closed set with connected complement such that $D\subseteq D^{\bullet}$. We give direct reconstruction methods (Theorems 3.6, 3.7, and 12.2) based on the monotonicity of the forward map $A\mapsto\Lambda(A)$. The main assumptions are self-adjointness of the coefficients, that $A_{0}$ satisfies a unique continuation principle, and that $A_{D}-A_{0}$ satisfies a local uniform definiteness condition inside $D$ near $\partial D^{\bullet}$; see Assumption 3.4 for the precise statement. In particular, inside $D$ we do not assume more regularity than $L^{\infty}$ for $A_{D}$, and $A_{D}-A_{0}$ can even be indefinite at a positive distance away from $\partial D^{\bullet}$.
As a consequence, for a given self-adjoint “background conductivity” $A_{0}$ satisfying the unique continuation principle, and any self-adjoint conductivities $A_{1}$ and $A_{2}$ (that equal $A_{0}$ near $\partial\Omega$), we have that
[…]
Hence, the obstructions from the anisotropic Calderón problem do not affect the inclusion detection problem under the definiteness condition. Practical reconstruction of inclusions is therefore sensible in a very high generality compared to reconstructing an anisotropic conductivity.
To provide some further intuition, consider a self-adjoint conductivity $A_{D}$, satisfying the definiteness condition, and with $\widetilde{A}_{D}$ defined from $A_{D}$ via (1.1) for a transformation $\Phi$. Let
$\widetilde{D}=\Phi(D)$
be the transformed inclusion under $\Phi$. If $\widetilde{D}^{\bullet}\neq D^{\bullet}$ then $\Phi$ will fail to be the identity in an open set intersecting $\partial D^{\bullet}$ (the part where it is deformed by $\Phi$), which also changes the background conductivity in $\widetilde{A}_{D}$ in that part of the domain via (1.1). We conclude for
$\widehat{D}=\mathop{\textup{supp}}(\widetilde{A}_{D}-A_{0}),$
if $\widehat{D}^{\bullet}\neq D^{\bullet}$, the definiteness condition cannot be satisfied inside $\widehat{D}$ near $\partial\widehat{D}^{\bullet}$. In particular, it may fail in part of the domain where the background conductivity in $\widetilde{A}_{D}$ is modified.
See also [8] for a stability result in a particular geometric setting. Moreover for Helmholtz scattering, there are uniqueness results for inclusion detection with some anisotropic coefficients [22, 6].
1.1. Outlining the main results
The reconstruction methods we present are based on monotonicity of the forward map via the following operator inequalities (Proposition 5.5):
$\int_{\Omega}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x\leq\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle\leq\int_{\Omega}A_{2}A_{1}^{-1}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x,$
where $\Lambda_{j}=\Lambda(A_{j})$ are the local ND maps, and $u_{2}$ is the electric potential for conductivity $A_{2}$ and Neumann condition $f$. The definiteness properties of $A_{2}A_{1}^{-1}(A_{2}-A_{1})$ turn out to be the same as for $A_{2}-A_{1}$ (Proposition 5.6).
The outer approach of the monotonicity method, given in our main results (Theorems 3.6, 3.7, and 12.2), constructs test-operators that use the monotonicity to check if a given admissible test-set $C$ fully contains $D^{\bullet}$ or not. Reconstruction is of the form (from Theorem 3.6):
$D\subseteq C\quad\text{\emph{if and only if}}\quad\Lambda_{C}^{-}\geq\Lambda(A_{D})\geq\Lambda_{C}^{+},$
for test-operators $\Lambda_{C}^{-}$ and $\Lambda_{C}^{+}$, and the local ND map $\Lambda(A_{D})$ for the unknown conductivity. In Theorem 3.7 the test-operators are linearized, which is beneficial for numerical implementation.
For Theorem 12.2 we also allow perfectly insulating and perfectly conducting parts in the unknown conductivity, and this is also used for the test-operators, which means that no conductivity bounds are needed for this approach. Theorem 12.2 is arguably the overall main result of the paper, although with the downside of requiring Lipschitz regularity of $\partial D^{\bullet}$ (which is unlike in Theorems 3.6 and 3.7, where no explicit boundary regularity is needed for the inclusions).
The paper is written in two parts, with Sections 210 dedicated to the finite perturbation setting, while Sections 1116 are dedicated to the extreme perturbation setting, following the approach for the isotropic case in [7, 16]. It should be noted that Theorem 12.2 cannot simply be obtained via a limit from Theorem 3.6. There is no short or straightforward proof for the difficult direction in the if and only if result, and it requires the introduction of several tools specific for extreme anisotropic coefficients; see [7, remark 5.4] for additional details in the isotropic case.
The proofs rely on the theory of localized potentials [21], for the existence of certain localizing solutions to the conductivity equation, based on a unique continuation principle. We generalize the result on localized potentials to anisotropic (even non-self-adjoint) coefficients in Theorem 6.3.
Under much stronger assumptions, we also give an inner approach of the monotonicity method in Theorems 4.2, 4.3, 12.4, and 12.5, which checks if a given open set $B$ is contained in $D^{\bullet}$.
1.2. Previous results for the monotonicity method
The monotonicity method is well-established for isotropic conductivities, and for other similar inverse coefficient problems. We mention some of the results here, with a focus on the Calderón problem. In [39] they use the monotonicity to give bounds on inclusions, and later [27, 30] prove that it is an exact reconstruction method for finite perturbations to the background conductivity, using the theory of localized potentials. It also includes the case with both positive and negative perturbations, if lower and upper bounds are known for the perturbed conductivity. This is expanded upon in [7], and now the perturbed conductivity can simultaneously have parts with finite positive and negative perturbations, as well as extreme parts that are perfectly insulating and perfectly conducting. The need for lower and upper bounds is also removed. Finally, in [19] the method is generalized to reconstruct collections of Lipschitz cracks.
The method has also been used in other specialized directions: degenerate and singular perturbations based on $A_{2}$-Muckenhoupt weights [16], inclusions in a fractional Calderón problem [28, 29], inclusions in a $p$-Laplace equation [5], for nonlinear materials [37], and for local uniqueness in an inverse coefficient problem for a non-resonant Schrödinger equation [26]. There are also rigorous connections to practically relevant electrode models [17, 18, 23, 31].
Certain reconstruction methods for the partial data isotropic Calderón problem also rely on the same type of methodology. For piecewise constant layered conductivities, there is an exact reconstruction method for both conductivity coefficient and the piecewise constant partition [12, 13]. In a finite-dimensional setting (where the partition is known), a general piecewise constant conductivity is reconstructed in [14], and likewise via a reformulation to a convex semidefinite optimization problem in [24].
1.3. Remarks on notation
All considered vector spaces are complex; this includes the classes of conductivity coefficients. We use the convention that inner products on complex Hilbert spaces are linear in the first entry and anti-linear in the second. We denote the Euclidean inner product as $z_{1}\cdot\overline{z_{2}}$ for $z_{1},z_{2}\in\mathbb{C}^{d}$, in particular the “dot” is bilinear. The Euclidean norm is denoted $\lvert\,\cdot\,\rvert$.
For self-adjoint operators $A,B\in\mathscr{L}(H)$ for a Hilbert space $H$, we write $A\geq B$ when ABA-B is positive semidefinite; this is the Loewner ordering of such operators. Moreover, for $A$ being a positive definite operator, $A^{1/2}$ denotes its unique positive definite square root.
We consistently make use of the essential support/supremum/infimum. We denote the interior of a set $C\subseteq\mathbb{R}^{d}$ as $C^{\circ}$. At certain occurrences (mainly from Section 11 and onwards) we denote by $K>0$ a generic constant that may change from line to line in the mathematical contents.
2. The forward problem
Let $\Omega$ be a bounded Lipschitz domain in $\mathbb{R}^{d}$, for $d\in\mathbb{N}\setminus\{1\}$, with connected complement. For a bounded matrix-valued function $A\in L^{\infty}(\Omega)^{d\times d}$, we use the norm
$\lVert A\rVert_{*}=\sup_{x\in\Omega}\lVert A(x)\rVert_{2},$
where $\lVert\,\cdot\,\rVert_{2}$ is the usual spectral norm for matrices, i.e. the operator norm on Euclidean space. Hence, we have
$\lvert A(x)\xi\rvert\leq\lVert A\rVert_{*}\lvert\xi\rvert$
for all $\xi\in\mathbb{C}^{d}$ and a.e. $x\in\Omega$.
We use the notation
$A^{\textup{R}}=\frac{1}{2}\bigl(A+A^{*}\bigr)\quad\text{and}\quad A^{\textup{I}}=\frac{1}{2\mathrm{i}}\bigl(A-A^{*}\bigr),$
such that
$A=A^{\textup{R}}+\mathrm{i}A^{\textup{I}},$
where we note that $A^{\textup{R}}$ and $A^{\textup{I}}$ are self-adjoint as matrices, i.e. Hermitian at every point in $\Omega$. This gives rise to the following general class of anisotropic coefficients:
$\mathcal{H}(\Omega)=\{\,A\in L^{\infty}(\Omega)^{d\times d}\mid\exists c>0\colon A^{\textup{R}}\geq cI\text{ in the Loewner order in $\Omega$}\,\}.$
The condition on $A$, the uniform positive definiteness of $A^{\textup{R}}$, can also explicitly be written:
$\inf_{x\in\Omega}\mathop{\textup{Re}}\bigl(A(x)\xi\cdot\overline{\xi}\bigr)=\inf_{x\in\Omega}A^{\textup{R}}(x)\xi\cdot\overline{\xi}\geq c\lvert\xi\rvert^{2}$
for all $\xi\in\mathbb{C}^{d}$; there are other classes of strongly elliptic coefficients (when restricting to $\xi\in\mathbb{R}^{d}$) that we do not consider here, see e.g. [35, chapter 4]. We also have the subset where $A$ is self-adjoint, i.e. when $A=A^{\textup{R}}$:
$\mathcal{H}_{\textup{SA}}(\Omega)=\{\,A\in\mathcal{H}(\Omega)\mid A^{\textup{I}}=0\,\}.$
Consider the partial data anisotropic conductivity problem for $A\in\mathcal{H}(\Omega)$:
[…]
Here $\nu$ is the outer unit normal to $\Omega$, $\Gamma\subseteq\partial\Omega$ is a non-empty open boundary piece, and the mean-free current density $f$ belongs to
$L^{2}_{\diamond}(\Gamma)=\{\,f\in L^{2}(\Gamma)\mid\langle f,1\rangle=0\,\}.$
We denote by $\langle\,\cdot\,$$\cdot\,\rangle$ the standard inner product on $L^{2}(\Gamma)$, and $\lVert\,\cdot\,\rVert$ is the associated norm. We likewise denote
$H^{1}_{\diamond}(\Omega)=\{\,u\in H^{1}(\Omega)\mid\langle u|_{\Gamma},1\rangle=0\,\}.$
A Poincaré inequality, related to the mean-free condition in $H^{1}_{\diamond}(\Omega)$, enables that
$\lVert v\rVert_{H^{1}_{\diamond}(\Omega)}=\Bigl(\int_{\Omega}\lvert\nabla v\rvert^{2}\,\mathrm{d}x\Bigr)^{1/2}$
becomes a norm on $H^{1}_{\diamond}(\Omega)$, equivalent with the usual $H^{1}$-norm. Moreover, the following continuity and coercivity estimates hold for all $w,v\in H^{1}_{\diamond}(\Omega)$:
$\bigl\lvert\int_{\Omega}A\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x\bigr\rvert$ $\leq\lVert A\rVert_{*}\lVert w\rVert_{H^{1}_{\diamond}(\Omega)}\lVert v\rVert_{H^{1}_{\diamond}(\Omega)},$
$\bigl\lvert\int_{\Omega}A\nabla v\cdot\overline{\nabla v}\,\mathrm{d}x\bigr\rvert$ $\geq\mathop{\textup{Re}}\int_{\Omega}A\nabla v\cdot\overline{\nabla v}\,\mathrm{d}x\geq c\lVert v\rVert_{H^{1}_{\diamond}(\Omega)}^{2}.$
By the Lax–Milgram lemma, there exists a unique solution $u=u_{f}^{A}\in H^{1}_{\diamond}(\Omega)$ to the corresponding weak problem associated to (2.1):
$\int_{\Omega}A\nabla u\cdot\overline{\nabla v}\,\mathrm{d}x=\langle f,v\rangle$ (2.2)
for all $v\in H^{1}_{\diamond}(\Omega)$. We let $\Lambda(A)\in\mathscr{L}(L^{2}_{\diamond}(\Gamma))$ denote the associated local ND map,
$\Lambda(A)f=u_{f}^{A}|_{\Gamma},$
and the forward problem is the nonlinear mapping $\Lambda\colon\mathcal{H}(\Omega)\to\mathscr{L}(L^{2}_{\diamond}(\Gamma))$.
From (2.2) it holds that for any $A_{1},A_{2}\in\mathcal{H}(\Omega)$ and $f,g\in L^{2}_{\diamond}(\Gamma)$, we have
$\langle f,\Lambda(A_{1})g\rangle=\int_{\Omega}A_{2}\nabla u_{f}^{A_{2}}\cdot\overline{\nabla u_{g}^{A_{1}}}\,\mathrm{d}x.$ (2.3)
Thus the adjoint operator satisfies
$\langle f,\Lambda(A_{1})^{*}g\rangle=\overline{\langle g,\Lambda(A_{1})f\rangle}=\int_{\Omega}\overline{A_{2}\nabla u_{g}^{A_{2}}}\cdot\nabla u_{f}^{A_{1}}\,\mathrm{d}x=\int_{\Omega}A_{2}^{*}\nabla u_{f}^{A_{1}}\cdot\overline{\nabla u_{g}^{A_{2}}}\,\mathrm{d}x,$
and therefore
$\langle f,\Lambda(A^{*})g\rangle=\int_{\Omega}A\nabla u_{f}^{A}\cdot\overline{\nabla u_{g}^{A^{*}}}\,\mathrm{d}x=\langle f,\Lambda(A)^{*}g\rangle.$
Specifically $\Lambda(A^{*})=\Lambda(A)^{*}$, meaning that $\Lambda(A)$ is self-adjoint if $A\in\mathcal{H}_{\textup{SA}}(\Omega)$.
Of particular interest are the quadratic forms
$\langle f,\Lambda(A)f\rangle=\int_{\Omega}A\nabla u_{f}^{A}\cdot\overline{\nabla u_{f}^{A}}\,\mathrm{d}x\quad\text{and}\quad\langle f,\Lambda(A)^{*}f\rangle=\int_{\Omega}A^{*}\nabla u_{f}^{A}\cdot\overline{\nabla u_{f}^{A}}\,\mathrm{d}x.$
We also write
$\Lambda^{\!\textup{R}}(A)=\frac{1}{2}\bigl(\Lambda(A)+\Lambda(A)^{*}\bigr)\quad\text{and}\quad\Lambda^{\!\textup{I}}(A)=\frac{1}{2\mathrm{i}}\bigl(\Lambda(A)-\Lambda(A)^{*}\bigr),$
so for the forward problem we have
$\Lambda=\Lambda^{\!\textup{R}}+\mathrm{i}\Lambda^{\!\textup{I}},$
where $\Lambda^{\!\textup{R}}(A)$ and $\Lambda^{\!\textup{I}}(A)$ are self-adjoint operators. In terms of the quadratic forms we thus have
$\langle f,\Lambda^{\!\textup{R}}(A)f\rangle$ $=\mathop{\textup{Re}}\langle f,\Lambda(A)f\rangle=\int_{\Omega}A^{\textup{R}}\nabla u_{f}^{A}\cdot\overline{\nabla u_{f}^{A}}\,\mathrm{d}x=\int_{\Omega}\lvert(A^{\textup{R}})^{1/2}\nabla u_{f}^{A}\rvert^{2}\,\mathrm{d}x,$ (2.4)
$\langle f,\Lambda^{\!\textup{I}}(A)f\rangle$ $=-\mathop{\textup{Im}}\langle f,\Lambda(A)f\rangle=-\int_{\Omega}A^{\textup{I}}\nabla u_{f}^{A}\cdot\overline{\nabla u_{f}^{A}}\,\mathrm{d}x.$ (2.5)
As one would expect, when multiplying with a complex scalar $\kappa\in\mathbb{C}$, we get
$(\kappa\Lambda)^{\textup{R}}=\kappa^{\textup{R}}\Lambda^{\!\textup{R}}-\kappa^{\textup{I}}\Lambda^{\!\textup{I}}\quad\text{and}\quad(\kappa\Lambda)^{\textup{I}}=\kappa^{\textup{I}}\Lambda^{\!\textup{R}}+\kappa^{\textup{R}}\Lambda^{\!\textup{I}}.$
3. Outer reconstruction of general anisotropic inclusions
In the following we will assume that a so-called “background conductivity” $A_{0}\in\mathcal{H}_{\textup{SA}}(\Omega)$ is known. For an unknown $A_{D}\in\mathcal{H}_{\textup{SA}}(\Omega)$, we call
$D=\mathop{\textup{supp}}(A_{D}-A_{0})$
the inclusions, noting that $D$ may have several connected components. Our reconstruction methods will relate to the so-called outer shape of $D$.
Definition 3.1.
The outer shape $D^{\bullet}$ of $D$ is the smallest closed set with connected complement such that $D\subseteq D^{\bullet}$.
One of the key assumptions we will need, required for the existence of certain localized solutions, is that the background conductivity $A_{0}$ satisfies a unique continuation principle.
Definition 3.2.
Let $U\subseteq\overline{\Omega}$ be relatively open and connected. $A\in\mathcal{H}(\Omega)$ satisfies the weak unique continuation principle (UCP) in $U$ for the conductivity equation if only the trivial solution of
[…]
can be identically zero in a non-empty open subset of $U$, and only the trivial solution has vanishing Cauchy data, $u$ and $\nu\cdot(A\nabla u)$, on a non-empty open part of $U\cap\partial\Omega$. If $U=\overline{\Omega}$ we just say that $A$ satisfies the UCP.
Remark 3.3.
It should be noted that the “UCP from Cauchy data” follows from the “UCP from open subsets”, so for simplicity we are grouping both concepts together in Definition 3.2.
In spatial dimension $d=2$, the UCP holds for general real-valued $A\in\mathcal{H}_{\textup{SA}}(\Omega)$ (and even in some non-self-adjoint cases), see [1].
In spatial dimension $d\geq 3$, the UCP holds for real-valued Lipschitz continuous coefficients $A\in\mathcal{H}_{\textup{SA}}(\Omega)$, see [20], and with counterexamples to the UCP for $\alpha$-Hölder continuous coefficients with $\alpha\in(0,1)$, see [36, 34].
We note that there also exist UCP results related to classes of complex coefficients, e.g. with principally normal symbol [32, chapter 28].
We define the class of admissible test-inclusions, used for the reconstruction methods, as
[…]
Assumption 3.4.
  1. (i)
    Assume there are known bounds on $A_{D}$, i.e. scalars $0<\alpha\leq\beta$ such that
    $\alpha I\leq A_{D}\leq\beta I$
    in the Loewner order in $\Omega$.
  2. (ii)
    Assume that $D\Subset\Omega$ and is the closure of an open set.
  3. (iii)
    For every $x\in\partial D^{\bullet}$ and every open neighborhood $W$ of $x$, assume there exists a relatively open connected set $V\subset D^{\bullet}\cap W$ that intersects $\partial D^{\bullet}$, satisfying either of two options:
    1. (a)
      $A_{D}-A_{0}$ is positive semidefinite in $V$, and there exists an open ball $B\subset V$ on which $A_{D}-A_{0}$ is uniformly positive definite.
    2. (b)
      $A_{D}-A_{0}$ is negative semidefinite in $V$, and there exists an open ball $B\subset V$ on which $A_{D}-A_{0}$ is uniformly negative definite.
  4. (iv)
    Assume that $A_{0}$ satisfies the UCP.
We give a few remarks on these assumptions.
Remark 3.5 (Remarks on Assumption 3.4).
  1. (a)
    In case $\Gamma=\partial\Omega$, then $C\Subset\Omega$ in the definition of $\mathcal{A}$ can be replaced by $C\subseteq\overline{\Omega}$, and $D\Subset\Omega$ in Assumption 3.4(ii) can be replaced by $D\subseteq\overline{\Omega}$. Also, here “connected complement” is in $\mathbb{R}^{d}$.
  2. (b)
    Note that Assumption 3.4(iii) only relates to what happens close to $\partial D^{\bullet}$, so further into $D$ we may have that $A_{D}-A_{0}$ is e.g. indefinite. Moreover, the definiteness condition is not exactly at $\partial D^{\bullet}$ but just close to $\partial D^{\bullet}$. This allows for both a jump condition or a continuous deviation from $A_{0}$; in the isotropic case this mimics the condition of a strict local increase/decrease from the background conductivity when entering $D$ from the outside [16].
For a measurable set $C$, we define the test-coefficients
$A^{-}_{C}=\begin{dcases}\alpha I&\text{in }C\\ A_{0}&\text{in }\Omega\setminus C\end{dcases}\quad\text{and}\quad A^{+}_{C}=\begin{dcases}\beta I&\text{in }C\\ A_{0}&\text{in }\Omega\setminus C,\end{dcases}$
and the corresponding test-operators
[…]
Theorem 3.6.
Under Assumption 3.4(i), for any measurable $C\subseteq\overline{\Omega}$ we have
$D\subseteq C\quad\text{implies}\quad\Lambda^{-}_{C}\geq\Lambda(A_{D})\geq\Lambda^{+}_{C}$.
Under all of Assumption 3.4, for any $C\in\mathcal{A}$ we have
$\Lambda^{-}_{C}\geq\Lambda(A_{D})\geq\Lambda^{+}_{C}\quad\text{implies}\quad D\subseteq C.$
Proof.
The proof is given in Section 7. ∎
We also introduce linearized versions of the test-operators, which are more suitable for numerical computations. Note that for $A\in\mathcal{H}_{\textup{SA}}(\Omega)$ and $B\in L^{\infty}(\Omega)^{d\times d}$, the Fréchet derivative of $\Lambda$ evaluated at $A$ and in direction $B$ is given by (c.f. [15]):
$\langle f,\textup{D}\mkern-1.5mu\Lambda(A;B)f\rangle=-\int_{\Omega}B^{*}\nabla u_{f}^{A}\cdot\overline{\nabla u_{f}^{A}}\,\mathrm{d}x.$
To shorten the notation in the following result, we introduce the test-operators
$\textup{D}\mkern-1.5mu\Lambda^{+}_{C}$ $=\textup{D}\mkern-1.5mu\Lambda\bigl(A_{0};(\beta I-A_{0})\chi_{C}\bigr),$
$\textup{D}\mkern-1.5mu\Lambda^{-}_{C}$ $=\textup{D}\mkern-1.5mu\Lambda\bigl(A_{0};(A_{0}-\tfrac{\beta^{2}}{\alpha}I)\chi_{C}\bigr),$
where $\chi_{C}$ is a characteristic function on the set $C$.
Theorem 3.7.
Pick $0<\alpha\leq\beta$ such that Assumption 3.4(i) holds for both $A_{D}$ and $A_{0}$. For any measurable $C\subseteq\overline{\Omega}$ we have
$D\subseteq C\quad\text{implies}\quad\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\geq\Lambda(A_{D})-\Lambda(A_{0})\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}$.
Under all of Assumption 3.4, for any $C\in\mathcal{A}$ we have
$\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\geq\Lambda(A_{D})-\Lambda(A_{0})\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\quad\text{implies}\quad D\subseteq C.$
Proof.
The proof is given in Section 8. ∎
Remark 3.8.
Under the given assumptions, Theorem 3.6 gives
$D^{\bullet}=\cap\,\{\,C\in\mathcal{A}\mid\Lambda^{-}_{C}\geq\Lambda(A_{D})\geq\Lambda^{+}_{C}\,\}$,
while Theorem 3.7 gives
$D^{\bullet}=\cap\,\{\,C\in\mathcal{A}\mid\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\geq\Lambda(A_{D})-\Lambda(A_{0})\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\,\}$.
Moreover, from the proofs we also conclude:
  • If in Assumption 3.4(iii) there is only positive definiteness near $\partial D^{\bullet}$, we only need to check the inequalities $\Lambda(A_{D})\geq\Lambda^{+}_{C}$ or $\Lambda(A_{D})-\Lambda(A_{0})\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}$.
  • If in Assumption 3.4(iii) there is only negative definiteness near $\partial D^{\bullet}$, we only need to check the inequalities $\Lambda^{-}_{C}\geq\Lambda(A_{D})$ or $\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\geq\Lambda(A_{D})-\Lambda(A_{0})$.
4. Inner reconstruction of uniformly definite anisotropic inclusions
Under much stronger definiteness assumptions (not just near $\partial D^{\bullet}$ as before), we can also use the inner approach of the monotonicity method. We focus on the linearized version, so that the UCP requirement is only related to the background conductivity $A_{0}$.
Assumption 4.1.
  1. (i)
    Assume that we have either:
    1. (a)
      Positive definite inclusions, $A_{D}-A_{0}\geq cI$ in the Loewner order in $D$, for known $c>0$.
    2. (b)
      Negative definite inclusions, $A_{D}-A_{0}\leq-cI$ in the Loewner order in $D$, for known $c>0$.
  2. (ii)
    Assume that $D\Subset\Omega$ and is the closure of an open set.
  3. (iii)
    Assume that $A_{0}$ satisfies the UCP.
Again, to shorten the notation in the following results, we introduce the test-operators
$\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}$ $=\textup{D}\mkern-1.5mu\Lambda\bigl(A_{0};c(\tfrac{\alpha}{\beta})^{2}\chi_{B}I\bigr),$
$\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}$ $=\textup{D}\mkern-1.5mu\Lambda\bigl(A_{0};-c\chi_{B}I\bigr),$
where $\chi_{B}$ is a characteristic function on the set $B$; note the differences compared to the previous section.
Theorem 4.2.
Assume that $A_{D}$ has positive definite inclusions in $D$, corresponding to case (a) of Assumption 4.1(i). Let $A_{0}\geq\alpha I$ and assume $A_{D}\leq\beta I$ in the Loewner order in $\Omega$, for $\alpha,\beta>0$. For any measurable $B\subseteq\Omega$ we have
$B\subseteq D\quad\text{implies}\quad\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}$.
Under all of Assumption 4.1, for any open set $B\subseteq\Omega$ we have
$\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}\quad\text{implies}\quad B\subset D^{\bullet}.$
Proof.
The proof is given in Section 9. ∎
Theorem 4.3.
Assume that $A_{D}$ has negative definite inclusions in $D$, corresponding to case (b) of Assumption 4.1(i). For any measurable $B\subseteq\Omega$ we have
$B\subseteq D\quad\text{implies}\quad\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\leq\Lambda(A_{D})-\Lambda(A_{0}).$
Under all of Assumption 4.1, for any open set $B\subseteq\Omega$ we have
$\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\leq\Lambda(A_{D})-\Lambda(A_{0})\quad\text{implies}\quad B\subset D^{\bullet}.$
Proof.
The proof is given in Section 10. ∎
Remark 4.4.
Under the given assumptions, for positive definite inclusions, Theorem 4.2 gives
$D^{\circ}\subseteq\cup\,\{\,B\subseteq\Omega\text{ open ball}\mid\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}\,\}\subset D^{\bullet},$
while for negative definite inclusions, Theorem 4.3 gives
$D^{\circ}\subseteq\cup\,\{\,B\subseteq\Omega\text{ open ball}\mid\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\leq\Lambda(A_{D})-\Lambda(A_{0})\,\}\subset D^{\bullet}.$
One can of course replace “open ball” by some other basis of open sets for the Euclidean topology.
Remark 4.5.
In the isotropic case, one can replace the factor $(\frac{\alpha}{\beta})^{2}$ in $\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}$ by the larger factor $\frac{\alpha}{\beta}$. This is because the same replacement holds in Proposition 5.6(ii) for isotropic coefficients, see e.g. [30, example 4.4].
5. Operator inequalities for Neumann-to-Dirichlet maps
Although we will only consider coefficients from $\mathcal{H}_{\textup{SA}}(\Omega)$ in the reconstruction methods, we still prove the following more general inequalities based on $\mathcal{H}(\Omega)$ for possible future reference. See also [25, Lemma 2.1] for the isotropic case with complex conductivity.
Theorem 5.1.
Let $A_{1},A_{2}\in\mathcal{H}(\Omega)$ and $f\in L^{2}_{\diamond}(\Gamma)$. For $j\in\{1,2\}$ we denote $\Lambda_{j}=\Lambda(A_{j})$ and $u_{j}=u_{f}^{A_{j}}$. Assume $\kappa\in\mathbb{C}$ is such that $\kappa A_{1}\in\mathcal{H}(\Omega)$. We also define
$\mathcal{B}_{\kappa}$ $=(\kappa A_{1})^{\textup{I}}[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{1})^{\textup{I}},$
$\mathcal{C}_{\kappa}$ $=(\kappa A_{2})^{\textup{I}}[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{2})^{\textup{I}}+\mathrm{i}\bigl[(\kappa A_{2})^{\textup{R}}[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{2})^{\textup{I}}-(\kappa A_{2})^{\textup{I}}[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{2})^{\textup{R}}\bigr].$
Then
$\langle f,[\overline{\kappa}(\Lambda_{1}-\Lambda_{2})]^{\textup{R}}f\rangle$ $\geq\int_{\Omega}\bigl[[\kappa(A_{2}-A_{1})]^{\textup{R}}-\mathcal{B}_{\kappa}\bigr]\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x,$ (5.1)
$\langle f,[\overline{\kappa}(\Lambda_{1}-\Lambda_{2})]^{\textup{R}}f\rangle$ $\leq\int_{\Omega}\bigl[(\kappa A_{2})^{\textup{R}}[(\kappa A_{1})^{\textup{R}}]^{-1}[\kappa(A_{2}-A_{1})]^{\textup{R}}+\mathcal{C}_{\kappa}\bigr]\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.$ (5.2)
Proof.
We start by proving (5.1). A computation reveals that
$(\kappa A_{1})[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{1})^{*}=(\kappa A_{1})^{\textup{R}}+\mathcal{B}_{\kappa},$
which is used below, combined with (2.3) and (2.4):
$0$ $\leq\int_{\Omega}\bigl\lvert[(\kappa A_{1})^{\textup{R}}]^{1/2}\nabla u_{1}-[(\kappa A_{1})^{\textup{R}}]^{-1/2}(\kappa A_{1})^{*}\nabla u_{2}\bigr\rvert^{2}\,\mathrm{d}x$
$=\mathop{\textup{Re}}\Bigl(\kappa\int_{\Omega}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x\Bigr)-2\mathop{\textup{Re}}\Bigl(\kappa\int_{\Omega}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x\Bigr)$
$\phantom{={}}+\int_{\Omega}\bigl[(\kappa A_{1})^{\textup{R}}+\mathcal{B}_{\kappa}\bigr]\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x$
$=\langle f,[\overline{\kappa}(\Lambda_{1}-\Lambda_{2})]^{\textup{R}}f\rangle+\int_{\Omega}\bigl[(\kappa A_{1})^{\textup{R}}-(\kappa A_{2})^{\textup{R}}+\mathcal{B}_{\kappa}\bigr]\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.$
Next we prove (5.2). A computation reveals that
$(\kappa A_{2})^{*}[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{2})$ $=(\kappa A_{2})^{\textup{R}}[(\kappa A_{1})^{\textup{R}}]^{-1}(\kappa A_{2})^{\textup{R}}+\mathcal{C}_{\kappa}$
$=(\kappa A_{2})^{\textup{R}}+(\kappa A_{2})^{\textup{R}}[(\kappa A_{1})^{\textup{R}}]^{-1}[\kappa(A_{2}-A_{1})]^{\textup{R}}+\mathcal{C}_{\kappa},$
which is used below, combined with (2.3) and (2.4):
$0$ $\leq\int_{\Omega}\bigl\lvert[(\kappa A_{1})^{\textup{R}}]^{1/2}\nabla u_{1}-[(\kappa A_{1})^{\textup{R}}]^{-1/2}(\kappa A_{2})\nabla u_{2}\bigr\rvert^{2}\,\mathrm{d}x$
$=\mathop{\textup{Re}}\Bigl(\kappa\int_{\Omega}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x\Bigr)-2\mathop{\textup{Re}}\Bigl(\kappa\int_{\Omega}A_{2}\nabla u_{2}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x\Bigr)$
$\phantom{={}}+\int_{\Omega}\bigl[(\kappa A_{2})^{\textup{R}}+(\kappa A_{2})^{\textup{R}}[(\kappa A_{1})^{\textup{R}}]^{-1}[\kappa(A_{2}-A_{1})]^{\textup{R}}+\mathcal{C}_{\kappa}\bigr]\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x$
$=\langle f,[\overline{\kappa}(\Lambda_{2}-\Lambda_{1})]^{\textup{R}}f\rangle+\int_{\Omega}\bigl[(\kappa A_{2})^{\textup{R}}[(\kappa A_{1})^{\textup{R}}]^{-1}[\kappa(A_{2}-A_{1})]^{\textup{R}}+\mathcal{C}_{\kappa}\bigr]\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.\qed$
Remark 5.2.
Note that in the isotropic case, the large square bracket in $\mathcal{C}_{\kappa}$ vanishes (after the imaginary unit). Also, if $\kappa A_{1}\in\mathcal{H}_{\textup{SA}}(\Omega)$ then $\mathcal{B}_{\kappa}=0$ and if $\kappa A_{2}\in\mathcal{H}_{\textup{SA}}(\Omega)$ then $\mathcal{C}_{\kappa}=0$.
We get some corollaries of Theorem 5.1, based on the choice of $\kappa$.
Corollary 5.3.
In the setting of Theorem 5.1 with $\kappa=1$, we have
$\langle f,(\Lambda^{\!\textup{R}}_{1}-\Lambda^{\!\textup{R}}_{2})f\rangle$ $\geq\int_{\Omega}\bigl[A^{\textup{R}}_{2}-A^{\textup{R}}_{1}-\mathcal{B}_{1}\bigr]\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x,$
$\langle f,(\Lambda^{\!\textup{R}}_{1}-\Lambda^{\!\textup{R}}_{2})f\rangle$ $\leq\int_{\Omega}\bigl[A^{\textup{R}}_{2}(A^{\textup{R}}_{1})^{-1}(A^{\textup{R}}_{2}-A^{\textup{R}}_{1})+\mathcal{C}_{1}\bigr]\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.$
Corollary 5.4.
In the setting of Theorem 5.1 with $\kappa=-\mathrm{i}$, provided that $A^{\textup{I}}_{1}\in\mathcal{H}_{\textup{SA}}(\Omega)$, we have
$\langle f,(\Lambda^{\!\textup{I}}_{2}-\Lambda^{\!\textup{I}}_{1})f\rangle$ $\geq\int_{\Omega}\bigl[A^{\textup{I}}_{2}-A^{\textup{I}}_{1}-\mathcal{B}_{-\mathrm{i}}\bigr]\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x,$
$\langle f,(\Lambda^{\!\textup{I}}_{2}-\Lambda^{\!\textup{I}}_{1})f\rangle$ $\leq\int_{\Omega}\bigl[A^{\textup{I}}_{2}(A^{\textup{I}}_{1})^{-1}(A^{\textup{I}}_{2}-A^{\textup{I}}_{1})+\mathcal{C}_{-\mathrm{i}}\bigr]\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.$
What we will need for the monotonicity method is the following, corresponding to the special case of Theorem 5.1 where $A_{1},A_{2}\in\mathcal{H}_{\textup{SA}}(\Omega)$ and $\kappa=1$. We also provide a slightly different proof that is arguably easier to read.
Proposition 5.5.
Let $A_{1},A_{2}\in\mathcal{H}_{\textup{SA}}(\Omega)$ and $f\in L^{2}_{\diamond}(\Gamma)$. For $j\in\{1,2\}$ we denote $\Lambda_{j}=\Lambda(A_{j})$ and $u_{j}=u_{f}^{A_{j}}$. Then
$\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle$ $\geq\int_{\Omega}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x,$ (5.3)
$\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle$ $\leq\int_{\Omega}A_{2}A_{1}^{-1}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.$ (5.4)
Proof.
We start by proving (5.3), which is by the computation below combined with (2.3),
$\int_{\Omega}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x$
$\hskip 42.67912pt\leq\int_{\Omega}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x+\int_{\Omega}\lvert A_{1}^{1/2}\nabla(u_{1}-u_{2})\rvert^{2}\,\mathrm{d}x$
$\hskip 42.67912pt=\int_{\Omega}A_{2}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x+\int_{\Omega}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x-2\mathop{\textup{Re}}\Bigl(\int_{\Omega}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x\Bigr)$
$\hskip 42.67912pt=\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle.$
Next we prove (5.4). Using that
$A_{2}A_{1}^{-1}(A_{2}-A_{1})=A_{2}A_{1}^{-1}A_{2}-A_{2}$
is self-adjoint, we get
$\int_{\Omega}A_{2}A_{1}^{-1}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x$
$\hskip 42.67912pt\geq\int_{\Omega}A_{2}A_{1}^{-1}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x-\int_{\Omega}\lvert A_{1}^{1/2}\nabla u_{1}-A_{1}^{-1/2}A_{2}\nabla u_{2}\rvert^{2}\,\mathrm{d}x$
$\hskip 42.67912pt=-\int_{\Omega}A_{2}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x-\int_{\Omega}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x+2\mathop{\textup{Re}}\Bigl(\int_{\Omega}A_{2}\nabla u_{2}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x\Bigr)$
$\hskip 42.67912pt=\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle.\qed$
Note that the first inequality in Proposition 5.5 implies monotonicity of the ND maps, since
$\int_{\Omega}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x\leq\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle\leq\int_{\Omega}(A_{2}-A_{1})\nabla u_{1}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x.$ (5.5)
It is not immediately clear how one would make use of the second inequality in Proposition 5.5, due to the matrix product in the front. What we need to consider is when $A_{2}-A_{1}$ is positive/negative (semi)definite, if we can also infer that $A_{2}A_{1}^{-1}(A_{2}-A_{1})$ has the same properties. This is the result of the following proposition.
Proposition 5.6.
Let $c\geq 0$ and $V\subseteq\Omega$. Let $A_{1},A_{2}\in\mathcal{H}_{\textup{SA}}(\Omega)$ with $A_{1}\leq\beta I$ and $A_{2}\geq\alpha I$ in $V$ for $\alpha,\beta>0$.
  1. (i)
    If $A_{2}-A_{1}\geq cI$ in $V$, then $A_{2}A_{1}^{-1}(A_{2}-A_{1})\geq cI$ in $V$.
  2. (ii)
    If $A_{2}-A_{1}\leq-cI$ in $V$, then $A_{2}A_{1}^{-1}(A_{2}-A_{1})\leq-c(\tfrac{\alpha}{\beta})^{2}I$ in $V$.
Proof.
Let $A_{2}-A_{1}\geq cI$ in the Loewner order in $V$ for some $c\geq 0$. Let $\xi\in\mathbb{C}^{d}$ and note that $A_{1}^{-1}$ is positive definite. Now we put the pieces together:
$A_{2}A_{1}^{-1}(A_{2}-A_{1})\xi\cdot\overline{\xi}$ $=(A_{2}-A_{1})A_{1}^{-1}(A_{2}-A_{1})\xi\cdot\overline{\xi}+(A_{2}-A_{1})\xi\cdot\overline{\xi}$
$=A_{1}^{-1}(A_{2}-A_{1})\xi\cdot\overline{(A_{2}-A_{1})\xi}+(A_{2}-A_{1})\xi\cdot\overline{\xi}$
$\geq c\lvert\xi\rvert^{2},$
which concludes the proof of (i).
For (ii), we first prove the assertion that if
$A_{1}-A_{2}\geq cI$ (5.6)
for some $c\geq 0$, then
$A_{2}^{-1}-A_{1}^{-1}\geq c\beta^{-2}I.$ (5.7)
The semidefinite case ($c=0$) of (5.7) is a simple consequence of $t\mapsto-t^{-1}$ being an operator monotone function, c.f. [9, chapter 7]. To prove (5.7), first note that we have
$A_{2}^{-1}-A_{1}^{-1}$ $=A_{1}^{-1}(A_{1}-A_{2})A_{2}^{-1}$
$=A_{1}^{-1}(A_{1}-A_{2})A_{1}^{-1}+A_{1}^{-1}(A_{1}-A_{2})(A_{2}^{-1}-A_{1}^{-1})$
$=A_{1}^{-1}\bigl[A_{1}-A_{2}+(A_{1}-A_{2})A_{2}^{-1}(A_{1}-A_{2})\bigr]A_{1}^{-1}.$
Note that $A_{1}^{-1}\geq\beta^{-1}I$ and likewise that $A_{2}^{-1}$ is positive definite. Hence, we have
$(A_{2}^{-1}-A_{1}^{-1})\xi\cdot\overline{\xi}$ $=(A_{1}-A_{2})A_{1}^{-1}\xi\cdot\overline{A_{1}^{-1}\xi}+A_{2}^{-1}(A_{1}-A_{2})A_{1}^{-1}\xi\cdot\overline{(A_{1}-A_{2})A_{1}^{-1}\xi}$
$\geq c\lvert A_{1}^{-1}\xi\rvert^{2}\geq c\beta^{-2}\lvert\xi\rvert^{2}.$
Now assume that in $V$ we have (5.6) for some $c\geq 0$. Then (5.7) implies
$A_{2}A_{1}^{-1}(A_{1}-A_{2})\xi\cdot\overline{\xi}=(A_{2}^{-1}-A_{1}^{-1})A_{2}\xi\cdot\overline{A_{2}\xi}\geq c\beta^{-2}\lvert A_{2}\xi\rvert^{2}\geq c(\tfrac{\alpha}{\beta})^{2}\lvert\xi\rvert^{2},$
which concludes the proof of (ii). ∎
6. Localized potentials for anisotropic coefficients
We extend localization results from [21] to anisotropic coefficients in $\mathcal{H}(\Omega)$. For a measurable set $V\subseteq\overline{\Omega}$ and $F\in L^{2}(V)^{d}$, we define $w=w_{F}^{A}$ as the unique solution in $H^{1}_{\diamond}(\Omega)$ to the variational problem
$\int_{\Omega}A^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x=\int_{V}F\cdot\overline{\nabla v}\,\mathrm{d}x$ (6.1)
for all $v\in H^{1}_{\diamond}(\Omega).$ We define the virtual measurement operator $L_{V}(A)\colon L^{2}(V)^{d}\to L^{2}_{\diamond}(\Gamma)$ as
$L_{V}(A)F=w_{F}^{A}|_{\Gamma}.$
In the following, $R$ is used to denote the range of an operator.
Proposition 6.1.
For $A\in\mathcal{H}(\Omega)$ and measurable sets $V,V_{1},V_{2}\subseteq\overline{\Omega}$, there are the following properties of the virtual measurement operator:
  1. (i)
    $R(L_{V}(A))$ is independent of the function values $A|_{V}$,
  2. (ii)
    $V_{1}\subseteq V_{2}$ implies $R(L_{V_{1}}(A))\subseteq R(L_{V_{2}}(A))$,
  3. (iii)
    The adjoint $L_{V}(A)^{*}\colon L^{2}_{\diamond}(\Gamma)\to L^{2}(V)^{d}$ is given by $L_{V}(A)^{*}f=\nabla u_{f}^{A}|_{V}$.
Proof.
(i): Let $A_{1},A_{2}\in\mathcal{H}(\Omega)$ such that $A_{1}=A_{2}$ outside $V$. We will prove that $R(L_{V}(A_{1}))=R(L_{V}(A_{2}))$. Let $\varphi\in R(L_{V}(A_{1}))$, i.e. $\varphi=w|_{\Gamma}$ with $w=w_{F_{1}}^{A_{1}}$ for some $F_{1}\in L^{2}(V)^{d}$. Now define $F_{2}=(A_{2}-A_{1})^{*}\nabla w|_{V}\in L^{2}(V)^{d}$. Then (6.1) gives
$\int_{\Omega}A_{2}^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x$ $=\int_{\Omega}(A_{2}-A_{1})^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x+\int_{\Omega}A_{1}^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x$
$=\int_{V}(A_{2}-A_{1})^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x+\int_{V}F_{1}\cdot\overline{\nabla v}\,\mathrm{d}x$
$=\int_{V}(F_{1}+F_{2})\cdot\overline{\nabla v}\,\mathrm{d}x$
for all $v\in H^{1}_{\diamond}(\Omega)$. Hence, we have that $\varphi=L_{V}(A_{2})(F_{1}+F_{2})$ so indeed $\varphi\in R(L_{V}(A_{2}))$. Repeating the above proof with $A_{1}$ and $A_{2}$ interchanged results in $R(L_{V}(A_{1}))=R(L_{V}(A_{2}))$ as desired.
(ii): Assume $V_{1}\subseteq V_{2}$, and let $\varphi=L_{V_{1}}(A)F=w|_{\Gamma}$ for some $F\in L^{2}(V_{1})^{d}$. Let $\widetilde{F}$ denote the extension by zero of $F$ on the set $V_{2}$. Then
$\int_{\Omega}A^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x=\int_{V_{1}}F\cdot\overline{\nabla v}\,\mathrm{d}x=\int_{V_{2}}\widetilde{F}\cdot\overline{\nabla v}\,\mathrm{d}x$
for all $v\in H^{1}_{\diamond}(\Omega)$. Hence, we also have $\varphi=L_{V_{2}}(A)\widetilde{F}$, i.e. $R(L_{V_{1}}(A))\subseteq R(L_{V_{2}}(A))$.
(iii): Let $F\in L^{2}(V)^{d}$ and $f\in L^{2}_{\diamond}(\Gamma)$, then from the variational forms (6.1) and (2.2), for $w=w_{F}^{A}$ and $u=u_{f}^{A}$, we obtain
$\langle L_{V}(A)^{*}f,F\rangle_{L^{2}(V)^{d}}$ $=\langle f,L_{V}(A)F\rangle_{L^{2}(\Gamma)}=\langle f,w\rangle_{L^{2}(\Gamma)}$
$=\int_{\Omega}A\nabla u\cdot\overline{\nabla w}\,\mathrm{d}x=\int_{V}\overline{F}\cdot\nabla u\,\mathrm{d}x=\langle\nabla u|_{V},F\rangle_{L^{2}(V)^{d}}.$
Since this holds for all $F\in L^{2}(V)^{d}$, we conclude that $L_{V}(A)^{*}f=\nabla u|_{V}$. ∎
We note that range inclusions imply norm inequalities for the adjoint operators.
Lemma 6.2 (Lemma 2.5 in [21]).
Let $H$, $K_{1}$, and $K_{2}$ be Hilbert spaces and let $T_{j}\in\mathscr{L}(K_{j},H)$ for $j\in\{1,2\}$. Then
$R(T_{1})\subseteq R(T_{2})\qquad\text{if and only if}\qquad\exists C>0,\,\forall x\in H\colon\lVert T_{1}^{*}x\rVert_{K_{1}}\leq C\lVert T_{2}^{*}x\rVert_{K_{2}}.$
Finally, we are able to prove the existence of localized potentials.
Theorem 6.3.
Let $U\subset\overline{\Omega}$ be a relatively open connected set that intersects $\Gamma$ and let $B\subset U$ be a non-empty open set. Assume $A\in\mathcal{H}(\Omega)$ where both $A$ and $A^{*}$ satisfy the UCP in $U$. Then there are sequences $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ and $(u_{j})$ in $H^{1}_{\diamond}(\Omega)$, with $u_{j}=u_{f_{j}}^{A}$, such that
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
Proof.
We let $V_{1}\Subset B\subset V\subseteq U$ where $V_{1}$ and $V$ are relatively open non-empty connected sets, and define $V_{2}=\Omega\setminus\overline{V}$. These sets are chosen such that $V$ intersects $\Gamma$ and that $\partial(\Omega\setminus\overline{V_{j}})$ are Lipschitz continuous.
We start by proving that $R(L_{V_{1}}(A))$ and $R(L_{V_{2}}(A))$ have a trivial intersection. Hence, let $\varphi\in R(L_{V_{1}}(A))\cap R(L_{V_{2}}(A))$, i.e. $\varphi=w_{j}|_{\Gamma}$ with $w_{j}=w_{F_{j}}^{A}$ for some $F_{j}\in L^{2}(V_{j})^{d}$ for $j\in\{1,2\}$. From (6.1) and the weak form of the conductivity problem (2.2), with conductivity $A^{*}$, we thus have
$-\nabla\cdot(A^{*}\nabla w_{j})$ $=0\text{ in }\Omega\setminus\overline{V_{j}}$
$\nu\cdot(A^{*}\nabla w_{j})$ $=0\text{ on }(\partial\Omega)\setminus\overline{V_{j}}.$
In particular $\nu\cdot(A^{*}\nabla w_{j})\equiv 0$ on $\partial V\cap\Gamma$, i.e. the Cauchy data of $w_{1}$ and $w_{2}$ coincide on $\partial V\cap\Gamma$. Unique continuation entails that $w_{1}=w_{2}$ in $V\setminus\overline{V_{1}}$. We may now “glue” $w_{1}$ and $w_{2}$ together, using that they agree in $V\setminus\overline{V_{1}}$:
$u=\begin{dcases}w_{1}&\text{in }\Omega\setminus\overline{V_{1}},\\ w_{2}&\text{in }V_{1}.\end{dcases}$
Now we have $\nabla\cdot(A^{*}\nabla u)=0$ in $\Omega$ and $\nu\cdot(A^{*}\nabla u)=\nu\cdot(A^{*}\nabla w_{1})\equiv 0$ on $\partial\Omega$. By uniqueness of such a solution in $H^{1}_{\diamond}(\Omega)$, it follows that $u\equiv 0$ in $\Omega$, so in particular $\varphi=w_{1}|_{\Gamma}=u|_{\Gamma}\equiv 0$, i.e.
$R(L_{V_{1}}(A))\cap R(L_{V_{2}}(A))=\{0\}.$
Next we prove that $R(L_{V_{1}}(A))$ is non-trivial. Suppose that $0\equiv L_{V_{1}}(A)^{*}f=\nabla u_{f}^{A}|_{V_{1}}$, then $u_{f}^{A}$ is a constant function in $V_{1}$, and also in $V$ by unique continuation. In particular, $f=\nu\cdot(A\nabla u_{f}^{A})|_{\Gamma}\equiv 0$ on $\partial V\cap\Gamma$. Thus, $L_{V_{1}}(A)^{*}$ is injective on
$W=\{\,f\in L^{2}_{\diamond}(\Gamma)\mid\mathop{\textup{supp}}f\subseteq\partial V\cap\Gamma\,\}.$
Since $\overline{R(L_{V_{1}}(A))}=N(L_{V_{1}}(A)^{*})^{\perp}$ then $R(L_{V_{1}}(A))$ contains a dense subset of $W$ and must be non-trivial.
In total we have that $R(L_{V_{1}}(A))\not\subseteq R(L_{V_{2}}(A))$. From Lemma 6.2 this implies the existence of a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A}$ satisfy (c.f. Proposition 6.1(iii)):
$\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x\geq\int_{V_{1}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\lVert L_{V_{1}}(A)^{*}f_{j}\rVert_{L^{2}(V_{1})^{d}}^{2}\to\infty$
and
$\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x\leq\int_{V_{2}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\lVert L_{V_{2}}(A)^{*}f_{j}\rVert_{L^{2}(V_{2})^{d}}^{2}\to 0$
for $j\to\infty$. ∎
The following result shows that the sequence $(f_{j})$ of Neumann conditions (e.g. from Theorem 6.3) such that the corresponding power $\lvert\nabla u_{j}\rvert^{2}$ is localized, can be chosen independently of the function values of $A$ on the set where the power tends to zero.
Proposition 6.4.
Let $B$ and $U$ be as in Theorem 6.3, and assume for $A_{1}\in\mathcal{H}(\Omega)$ and $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ that $u_{j}=u_{f_{j}}^{A_{1}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
If $A_{2}\in\mathcal{H}(\Omega)$ with $\mathop{\textup{supp}}(A_{1}-A_{2})\subseteq\overline{\Omega}\setminus U$, then $\widehat{u}_{j}=u_{f_{j}}^{A_{2}}$ also satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x=0.$
Proof.
Let $V_{1}=B$ and $V_{2}=\overline{\Omega}\setminus U$. From Proposition 6.1(iii) it amounts to proving
$\lim_{j\to\infty}\lVert L_{V_{1}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{1})^{d}}=\infty\quad\text{and}\quad\lim_{j\to\infty}\lVert L_{V_{2}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{2})^{d}}=0.$´
From Proposition 6.1(i) we have that $R(L_{V_{2}}(A_{1}))=R(L_{V_{2}}(A_{2}))$. Hence a combination of Proposition 6.1(iii) and Lemma 6.2 gives
$\lVert L_{V_{2}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{2})^{d}}\leq C\lVert L_{V_{2}}(A_{1})^{*}f_{j}\rVert_{L^{2}(V_{2})^{d}}\to 0\text{ for }j\to\infty.$
By the same argumentation, we also have $R(L_{V_{1}\cup V_{2}}(A_{1}))=R(L_{V_{1}\cup V_{2}}(A_{2}))$, which implies
$\lVert L_{V_{1}\cup V_{2}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{1}\cup V_{2})^{d}}$ $\geq c\lVert L_{V_{1}\cup V_{2}}(A_{1})^{*}f_{j}\rVert_{L^{2}(V_{1}\cup V_{2})^{d}}$
$\geq c\lVert L_{V_{1}}(A_{1})^{*}f_{j}\rVert_{L^{2}(V_{1})^{d}}\to\infty\text{ for }j\to\infty.$
Finally we combine the two above limits, using Proposition 6.1(iii),
$\lVert L_{V_{1}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{1})^{d}}^{2}=\lVert L_{V_{1}\cup V_{2}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{1}\cup V_{2})^{d}}^{2}-\lVert L_{V_{2}}(A_{2})^{*}f_{j}\rVert_{L^{2}(V_{2})^{d}}^{2}\to\infty\text{ for }j\to\infty,$
thereby concluding the proof. ∎
7. Proof of Theorem 3.6
The proof that $D\subseteq C$ implies $\Lambda^{-}_{C}\geq\Lambda(A_{D})\geq\Lambda^{+}_{C}$ follows directly from Assumption 3.4(i) and (5.5), since we have
$A^{-}_{C}\leq A_{D}\leq A^{+}_{C}$
in the Loewner order in $\Omega$.
Now assume that $D\not\subseteq C$, i.e. that $D^{\bullet}\not\subseteq C$. Since both sets are closures of opens sets and have connected complements, there is a relatively open set that connects $D^{\bullet}\setminus C$ with $\Gamma$ that we may use for localization. Specifically, there exist a relatively open connected set $U\subset\overline{\Omega}\setminus C$ that intersects $\Gamma$ and an open ball $B\subset U\cap D$. Moreover, Assumption 3.4(iii) ensures that we can pick $B$ and $U$ such that we arrive at one of two cases:
  • Case 1: $A_{D}-A_{0}$ is positive semidefinite in $U$ and uniformly positive definite in $B$.
  • Case 2: $A_{D}-A_{0}$ is negative semidefinite in $U$ and uniformly negative definite in $B$.
In the first case we will prove that $\Lambda(A_{D})\not\geq\Lambda^{+}_{C}$, and in the second case we will prove that $\Lambda^{-}_{C}\not\geq\Lambda(A_{D})$. Together the two cases prove the contrapositive formulation of the assertion
$\Lambda^{-}_{C}\geq\Lambda(A_{D})\geq\Lambda^{+}_{C}\quad\text{\emph{implies}}\quad D\subseteq C.$
Case 1
Using Theorem 6.3 and Proposition 6.4 (as $C\subset\Omega\setminus U$), there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A^{+}_{C}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
Note that $A_{C}^{+}=A_{0}$ in $U$ where $A_{D}-A_{0}$ is positive semidefinite, and also in $B$ where $A_{D}-A_{0}$ is uniformly positive definite. From Proposition 5.6(ii) then $A_{0}A_{D}^{-1}(A_{0}-A_{D})$ is negative semidefinite in $U$, and there exists a scalar $\tilde{c}>0$ such that
$A_{0}A_{D}^{-1}(A_{0}-A_{D})\leq-\tilde{c}I\text{ in }B.$
Now Proposition 5.5 gives
$\langle f_{j},\bigl[\Lambda(A_{D})-\Lambda^{+}_{C}\bigr]f_{j}\rangle$ $\leq\int_{\Omega}A^{+}_{C}A_{D}^{-1}(A^{+}_{C}-A_{D})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x$
$\leq\lVert A^{+}_{C}A_{D}^{-1}(A^{+}_{C}-A_{D})\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x-\tilde{c}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x.$
We conclude that
$\lim_{j\to\infty}\langle f_{j},\bigl[\Lambda(A_{D})-\Lambda^{+}_{C}\bigr]f_{j}\rangle=-\infty,$
in particular, $\Lambda(A_{D})\not\geq\Lambda^{+}_{C}$.
Case 2
Using Theorem 6.3 and Proposition 6.4 (as $C\subset\Omega\setminus U$), there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A^{-}_{C}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
Note that $A_{C}^{-}=A_{0}$ in $U$ where $A_{D}-A_{0}$ is negative semidefinite, and also in $B$ where $A_{D}-A_{0}$ is uniformly negative definite. From Proposition 5.5 we have
$\langle f_{j},\bigl[\Lambda^{-}_{C}-\Lambda(A_{D})\bigr]f_{j}\rangle$ $\leq\int_{\Omega}(A_{D}-A^{-}_{C})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x$
$\leq\lVert A_{D}-A^{-}_{C}\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x-c\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x$
for a scalar $c>0$. We conclude that
$\lim_{j\to\infty}\langle f_{j},\bigl[\Lambda^{-}_{C}-\Lambda(A_{D})\bigr]f_{j}\rangle=-\infty,$
in particular, $\Lambda^{-}_{C}\not\geq\Lambda(A_{D})$.
8. Proof of Theorem 3.7
Assume $D\subseteq C$ and let $u_{0}=u_{f}^{A_{0}}$. From Proposition 5.5 we have
$\langle f,\bigl[\Lambda(A_{D})-\Lambda(A_{0})-\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\bigr]f\rangle$ $\geq\int_{\Omega}(A_{0}-A_{D})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x+\int_{C}(\beta I-A_{0})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x$
$=\int_{D}(A_{0}-A_{D})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x+\int_{C}(\beta I-A_{0})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x$
$\geq\int_{C\setminus D}(\beta I-A_{0})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x\geq 0.$
For the next part we need an additional bound. Let $\xi\in\mathbb{C}^{d}$, then we have
$A_{0}A_{D}^{-1}A_{0}\xi\cdot\overline{\xi}=A_{D}^{-1}A_{0}\xi\cdot\overline{A_{0}\xi}\leq\alpha^{-1}\lvert A_{0}\xi\rvert^{2}\leq\tfrac{\beta^{2}}{\alpha}\lvert\xi\rvert^{2}.$ (8.1)
Note that $A_{0}\leq\beta I\leq\frac{\beta^{2}}{\alpha}I$. Now from Proposition 5.5 and (8.1) we get
$\langle f,\bigl[\Lambda(A_{0})-\Lambda(A_{D})+\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\bigr]f\rangle$
$\hskip-85.35826pt\geq\int_{\Omega}A_{0}A_{D}^{-1}(A_{D}-A_{0})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x+\int_{C}\bigl(\tfrac{\beta^{2}}{\alpha}I-A_{0}\bigr)\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x$
$\hskip-85.35826pt=\int_{D}(A_{0}-A_{0}A_{D}^{-1}A_{0})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x+\int_{C}\bigl(\tfrac{\beta^{2}}{\alpha}I-A_{0}\bigr)\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x$
$\hskip-85.35826pt\geq\int_{C\setminus D}\bigl(\tfrac{\beta^{2}}{\alpha}I-A_{0}\bigr)\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x\geq 0.$
This concludes the proof for the first part of the theorem statement.
Now assume that $D\not\subseteq C$, i.e. that $D^{\bullet}\not\subseteq C$. Since both sets are closures of opens sets and have connected complements, there is a relatively open set that connects $D^{\bullet}\setminus C$ with $\Gamma$ that we may use for localization. Specifically, there exist a relatively open connected set $U\subset\overline{\Omega}\setminus C$ that intersects $\Gamma$ and an open ball $B\subset U\cap D$. Moreover, Assumption 3.4(iii) ensures that we can pick $B$ and $U$ such that we arrive at one of two cases:
  • Case 1: $A_{D}-A_{0}$ is positive semidefinite in $U$ and uniformly positive definite in $B$.
  • Case 2: $A_{D}-A_{0}$ is negative semidefinite in $U$ and uniformly negative definite in $B$.
In the first case we will prove that $\Lambda(A_{D})-\Lambda(A_{0})\not\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}$, and in the second case we will prove that $\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\not\geq\Lambda(A_{D})-\Lambda(A_{0})$. Together the two cases prove the contrapositive formulation of the assertion
$\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\geq\Lambda(A_{D})-\Lambda(A_{0})\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\quad\text{\emph{implies}}\quad D\subseteq C.$
For both cases we use the same sequence of localized potentials. From Theorem 6.3 there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A_{0}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
Case 1
We now combine Proposition 5.5 with Proposition 5.6(ii), using the facts that $A_{D}-A_{0}$ is positive semidefinite in $U$ and $A_{D}-A_{0}\geq cI$ in $B$ for some scalar $c>0$, which results in
$\langle f_{j},\bigl[\Lambda(A_{D})-\Lambda(A_{0})-\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\bigr]f_{j}\rangle$
$\hskip-113.81102pt\leq\int_{\Omega}A_{0}A_{D}^{-1}(A_{0}-A_{D})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x+\int_{C}(\beta I-A_{0})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x$
$\hskip-113.81102pt\leq\lVert A_{0}A_{D}^{-1}(A_{0}-A_{D})\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x-c\bigl(\tfrac{\alpha}{\beta}\bigr)^{2}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x+\lVert\beta I-A_{0}\rVert_{*}\int_{C}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x.$
Since $C\subset\Omega\setminus U$, we conclude that
$\lim_{j\to\infty}\langle f_{j},\bigl[\Lambda(A_{D})-\Lambda(A_{0})-\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\bigr]f_{j}\rangle=-\infty,$
in particular, $\Lambda(A_{D})-\Lambda(A_{0})\not\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}$.
Case 2
We use Proposition 5.5 together with the facts that $A_{D}-A_{0}$ is negative semidefinite in $U$ and $A_{D}-A_{0}\leq-cI$ in $B$ for some scalar $c>0$, which results in
$\langle f_{j},\bigl[\Lambda(A_{0})-\Lambda(A_{D})+\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\bigr]f_{j}\rangle$
$\hskip-113.81102pt\leq\int_{\Omega}(A_{D}-A_{0})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x+\int_{C}\bigl(\tfrac{\beta^{2}}{\alpha}I-A_{0}\bigr)\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x$
$\hskip-113.81102pt\leq\lVert A_{D}-A_{0}\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x-c\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x+\lVert\tfrac{\beta^{2}}{\alpha}I-A_{0}\rVert_{*}\int_{C}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x.$
Since $C\subset\Omega\setminus U$, we conclude that
$\lim_{j\to\infty}\langle f_{j},\bigl[\Lambda(A_{0})-\Lambda(A_{D})+\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\bigr]f_{j}\rangle=-\infty,$
in particular, $\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\not\geq\Lambda(A_{D})-\Lambda(A_{0})$.
9. Proof of Theorem 4.2
Assume $B\subseteq D$ and let $u_{0}=u_{f}^{A_{0}}$. By assumption, $A_{D}-A_{0}\geq cI$ in $D$ for some $c>0$, from which Proposition 5.6(ii) implies that
$A_{0}A_{D}^{-1}(A_{D}-A_{0})\geq c\bigl(\tfrac{\alpha}{\beta}\bigr)^{2}I\text{ in }D.$
By Proposition 5.5, we have
$\langle f,\bigl[\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}+\Lambda(A_{0})-\Lambda(A_{D})\bigr]f\rangle$ $\geq-c\bigl(\tfrac{\alpha}{\beta}\bigr)^{2}\int_{B}\lvert\nabla u_{0}\rvert^{2}\,\mathrm{d}x+\int_{\Omega}A_{0}A_{D}^{-1}(A_{D}-A_{0})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x$
$\geq c\bigl(\tfrac{\alpha}{\beta}\bigr)^{2}\int_{D\setminus B}\lvert\nabla u_{0}\rvert^{2}\,\mathrm{d}x\geq 0.$
Next, assume $B\not\subset D^{\bullet}$. Since $D^{\bullet}$ is closed and has connected complement, while $B$ is a non-empty open set, there exist a relatively open and connected set $U\subseteq\overline{\Omega}\setminus D^{\bullet}$ intersecting $\Gamma$ and $B$, and an open ball $\widehat{B}\subset U\cap B$. From Theorem 6.3 there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A_{0}}$ satisfy
$\lim_{j\to\infty}\int_{\widehat{B}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
Since $D\subset\Omega\setminus U$, Proposition 5.5 gives
$\langle f_{j},\bigl[\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}+\Lambda(A_{0})-\Lambda(A_{D})\bigr]f_{j}\rangle$ $\leq-c\bigl(\tfrac{\alpha}{\beta}\bigr)^{2}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x+\int_{\Omega}(A_{D}-A_{0})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x$
$\leq-c\bigl(\tfrac{\alpha}{\beta}\bigr)^{2}\int_{\widehat{B}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x+\lVert A_{D}-A_{0}\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x.$
We conclude that
$\lim_{j\to\infty}\langle f_{j},\bigl[\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}+\Lambda(A_{0})-\Lambda(A_{D})\bigr]f_{j}\rangle=-\infty,$
in particular, $\Lambda(A_{D})-\Lambda(A_{0})\not\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}$. This proves the contrapositive formulation of the assertion
$\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}\quad\text{\emph{implies}}\quad B\subset D^{\bullet}.$
10. Proof of Theorem 4.3
Assume $B\subseteq D$ and let $u_{0}=u_{f}^{A_{0}}$. By assumption, $A_{0}-A_{D}\geq cI$ in $D$, so Proposition 5.5 yields
$\langle f,\bigl[\Lambda(A_{D})-\Lambda(A_{0})-\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\bigr]f\rangle$ $\geq-c\int_{B}\lvert\nabla u_{0}\rvert^{2}\,\mathrm{d}x+\int_{\Omega}(A_{0}-A_{D})\nabla u_{0}\cdot\overline{\nabla u_{0}}\,\mathrm{d}x$
$\geq c\int_{D\setminus B}\lvert\nabla u_{0}\rvert^{2}\,\mathrm{d}x\geq 0.$
Next, assume $B\not\subset D^{\bullet}$. Since $D^{\bullet}$ is closed and has connected complement, while $B$ is a non-empty open set, there exist a relatively open and connected set $U\subseteq\overline{\Omega}\setminus D^{\bullet}$ intersecting $\Gamma$ and $B$, and an open ball $\widehat{B}\subset U\cap B$. From Theorem 6.3 there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A_{0}}$ satisfy
$\lim_{j\to\infty}\int_{\widehat{B}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
Since $D\subset\Omega\setminus U$, Proposition 5.5 gives
$\langle f_{j},\bigl[\Lambda(A_{D})-\Lambda(A_{0})-\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\bigr]f_{j}\rangle$ $\leq-c\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x+\int_{\Omega}A_{0}A_{D}^{-1}(A_{0}-A_{D})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x$
$\leq-c\int_{\widehat{B}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x+\lVert A_{0}A_{D}^{-1}(A_{0}-A_{D})\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x.$
We conclude that
$\lim_{j\to\infty}\langle f_{j},\bigl[\Lambda(A_{D})-\Lambda(A_{0})-\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\bigr]f_{j}\rangle=-\infty,$
in particular, $\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\not\leq\Lambda(A_{D})-\Lambda(A_{0})$. This proves the contrapositive formulation of the assertion
$\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\leq\Lambda(A_{D})-\Lambda(A_{0})\quad\text{\emph{implies}}\quad B\subset D^{\bullet}.$
11. The forward problem with extreme coefficients
We will now consider extreme inclusions, meaning that part of the domain $C_{0}$ is occupied by a perfectly insulating conductivity (vanishing conductivity) and part of the domain $C_{\infty}$ is occupied by a perfectly conducting conductivity (infinite conductivity). The extreme parts are inherently isotropic. We may formally denote such a coefficient as
$A=\begin{dcases}A_{\textup{F}}&\text{in }\Omega\setminus C\\ 0I&\text{in }C_{0}\\ \infty I&\text{in }C_{\infty},\end{dcases}$ (11.1)
where $C=C_{0}\cup C_{\infty}$ and where $A_{\textup{F}}\in\mathcal{H}(\Omega)$ refers to the finite/non-extreme part of the coefficient. For notational convenience, since we later define extensions into $C_{0}$ and an inner product related to $\Omega\setminus C_{0}$, we will always consider $A_{\textup{F}}$ to be defined in all of $\Omega$. We apply the following assumption in connection with such extreme inclusions.
Assumption 11.1.
Let $C_{0},C_{\infty}\Subset\Omega$ satisfy:
  • $C_{0}$ and $C_{\infty}$ are closures of open sets with Lipschitz boundaries.
  • $C_{0}\cap C_{\infty}=\emptyset$.
  • $\Omega\setminus C_{0}$ is connected.
It is worth noting that $C_{0}$ and $C_{\infty}$ are allowed to comprise several connected components, and are also allowed to be empty sets. These extreme parts can be introduced via certain boundary conditions:
$-\nabla\cdot(A_{\textup{F}}\nabla u)$ $=0\text{ in }\Omega\setminus C$
$\nu\cdot(A_{\textup{F}}\nabla u)$ $=\begin{dcases}f&\text{on }\Gamma\\ 0&\text{on }\partial(\Omega\setminus C_{0})\setminus\Gamma\end{dcases}$
$\nabla u$ $=0\text{ in }C_{\infty}^{\circ}$
$\int_{\partial C_{i}}\nu\cdot(A_{\textup{F}}\nabla u)\,\mathrm{d}S$ $=0\text{ for each component $C_{i}$ of $C_{\infty}$}.$
In the last condition the trace is taken from the exterior of $C_{\infty}$. It is worth noting here, that $u$ is only defined in $\Omega\setminus C_{0}$ from this PDE problem. A consistent choice of extending $u$ into $C_{0}$, which we will use throughout this paper, is the $H^{1}$-extension $Eu$ of $u$ defined below.
Definition 11.2.
Define $E\colon H^{1}(\Omega\setminus C_{0})\to H^{1}(\Omega)$ as
$Ew=\begin{dcases}w&\text{in }\Omega\setminus C_{0}\\ \widetilde{w}&\text{in }C_{0},\end{dcases}$
with $\widetilde{w}$ being the unique solution in $H^{1}(C_{0}^{\circ})$ of the Dirichlet problem
[…]
$\widetilde{w}$ $=w\text{ on }\partial C_{0}.$
It is clear that the extension $E$ depends on $C_{0}$ and $A_{\textup{F}}$, however to avoid introducing excessive notation, whenever we write $Eu$ it will be implicitly understood that we use the extension related to the $C_{0}$ and $A_{\textup{F}}$ from the definition of the specific electric potential $u$.
Next we consider the weak formulation in connection with extreme inclusions, and for this purpose we introduce the space
$\mathscr{H}_{C_{0}}^{C_{\infty}}=\{\,v\in H_{\diamond}^{1}(\Omega\setminus C_{0})\mid\nabla v=0\text{ in }C_{\infty}^{\circ}\,\}.$
The weak problem now becomes
$\int_{\Omega\setminus C_{0}}A_{\textup{F}}\nabla u\cdot\overline{\nabla v}\,\mathrm{d}x=\langle f,v\rangle$ (11.2)
for all $v\in\mathscr{H}_{C_{0}}^{C_{\infty}}$. Once again, the Lax–Milgram lemma ensures the existence of a unique solution $u=u_{f}^{A}\in\mathscr{H}_{C_{0}}^{C_{\infty}}$, and we have the local ND map $\Lambda(A)\colon f\mapsto u_{f}^{A}|_{\Gamma}$.
From (11.2), for any $f,g\in L^{2}_{\diamond}(\Gamma)$ and for any two pairs of extreme coefficients $A_{1}$ and $A_{2}$, with extreme inclusions $C=C_{0}\cup C_{\infty}$ for $A_{1}$ and where the perfectly conducting parts of $A_{2}$ are a subset of $C_{\infty}$, we have:
$\langle f,\Lambda(A_{1})g\rangle=\int_{\Omega\setminus C}A_{2}\nabla u_{f}^{A_{2}}\cdot\overline{\nabla Eu_{g}^{A_{1}}}\,\mathrm{d}x.$ (11.3)
In particular
$\langle f,\Lambda(A)g\rangle=\int_{\Omega\setminus C}A_{\textup{F}}\nabla u_{f}^{A}\cdot\overline{\nabla u_{g}^{A}}\,\mathrm{d}x=\int_{\Omega\setminus C_{0}}A_{\textup{F}}\nabla u_{f}^{A}\cdot\overline{\nabla u_{g}^{A}}\,\mathrm{d}x,$ (11.4)
for which $\Lambda(A)$ is self-adjoint if $A_{\textup{F}}\in\mathcal{H}_{\textup{SA}}(\Omega)$.
We moreover note that when $A_{\textup{F}}\in\mathcal{H}_{\textup{SA}}(\Omega)$, then
$\langle w,v\rangle_{A}=\int_{\Omega\setminus C_{0}}A_{\textup{F}}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x$ (11.5)
becomes an inner product on $\mathscr{H}_{C_{0}}^{\emptyset}\supseteq\mathscr{H}_{C_{0}}^{C_{\infty}}$, whose norm $\lVert\,\cdot\,\rVert_{A}$ is equivalent with the norm on $H^{1}_{\diamond}(\Omega\setminus C_{0})$. We will often make use of the norm equivalences between $\lVert\,\cdot\,\rVert_{H^{1}(\Omega\setminus C_{0})}$, $\lVert\nabla(\cdot)\rVert_{L^{2}(\Omega\setminus C_{0})^{d}}$, and $\lVert\,\cdot\,\rVert_{A}$ that hold on $\mathscr{H}_{C_{0}}^{\emptyset}$.
12. Reconstruction of general anisotropic and extreme inclusions
We now give the main result on reconstructing anisotropic inclusions also including extreme inclusions. We moreover avoid requiring conductivity bounds such as those needed for Theorem 3.6. The main disadvantages, however, are twofold:
  • We assume Lipschitz regularity of the outer inclusion boundaries, which is unlike in Theorems 3.6 and 3.7 where no explicit boundary regularity is needed for the inclusions.
  • For numerical implementation, a linearized version of the method such as Theorem 3.7 is very beneficial. For general inclusions this cannot be obtained in the extreme case. However, in the less general inner approach, one can still obtain linearized versions of the method for either perfectly insulating or perfectly conducting inclusions (Theorem 12.4).
We will assume to have a known (non-extreme) background conductivity $A_{0}\in\mathcal{H}_{\textup{SA}}(\Omega)$ and an unknown conductivity $A_{D}$ of the form
[…]
where $D_{0}$ and $D_{\infty}$ satisfy Assumption 11.1 (in place of $C_{0}$ and $C_{\infty}$), and with the finite part $A_{\textup{F}}\in\mathcal{H}_{\textup{SA}}(\Omega)$.
Once again, we define the inclusions as
$D=\mathop{\textup{supp}}(A_{D}-A_{0}),$
and proceed to determine the outer shape $D^{\bullet}$. We denote by
$D_{\textup{F}}=D\setminus(D_{0}\cup D_{\infty})$
the part of the domain occupied by non-extreme deviations from $A_{0}$.
We define the class of admissible test-inclusions, which in the extreme case also have Lipschitz regular boundary:
$\widehat{\mathcal{A}}$ $=\{\,C\Subset\Omega\mid C\text{ is the closure of an open set,}$
has connected complement,
[…]
Assumption 12.1.
  1. (i)
    Assume that $D\Subset\Omega$ and is the closure of an open set, and that $\partial D^{\bullet}$ is Lipschitz regular.
  2. (ii)
    For every $x\in\partial D^{\bullet}$ and every open neighborhood $W$ of $x$, assume there exists a relatively open connected set $V\subset D^{\bullet}\cap W$ that intersects $\partial D^{\bullet}$, satisfying either that $V\subseteq D_{0}$, $V\subseteq D_{\infty}$, or $V\subseteq D_{\textup{F}}$. In the latter case, $V$ can be picked such that either of two options hold:
    1. (a)
      $A_{\textup{F}}-A_{0}$ is positive semidefinite in $V$, and there exists an open ball $B\subset V$ on which $A_{\textup{F}}-A_{0}$ is uniformly positive definite.
    2. (b)
      $A_{\textup{F}}-A_{0}$ is negative semidefinite in $V$, and there exists an open ball $B\subset V$ on which $A_{\textup{F}}-A_{0}$ is uniformly negative definite.
  3. (iii)
    Assume that $A_{0}$ satisfies the UCP.
For $C\in\widehat{\mathcal{A}}$, we define the test-coefficients
$A_{C}^{\emptyset}=\begin{dcases}0I&\text{in }C\\ A_{0}&\text{in }\Omega\setminus C\end{dcases}\quad\text{and}\quad A_{\emptyset}^{C}=\begin{dcases}\infty I&\text{in }C\\ A_{0}&\text{in }\Omega\setminus C,\end{dcases}$
and the corresponding test-operators
$\Lambda_{C}^{\emptyset}=\Lambda(A_{C}^{\emptyset})\quad\text{and}\quad\Lambda_{\emptyset}^{C}=\Lambda(A_{\emptyset}^{C}).$
Theorem 12.2.
For any $C\in\widehat{\mathcal{A}}$ we have
$D\subseteq C\quad\text{implies}\quad\Lambda_{C}^{\emptyset}\geq\Lambda(A_{D})\geq\Lambda_{\emptyset}^{C}.$
Under Assumption 12.1, for any $C\in\widehat{\mathcal{A}}$ we have
$\Lambda_{C}^{\emptyset}\geq\Lambda(A_{D})\geq\Lambda_{\emptyset}^{C}\quad\text{implies}\quad D\subseteq C.$
Proof.
The proof is given in Section 16. ∎
Remark 12.3.
Under Assumptions 12.1, Theorem 12.2 gives
$D^{\bullet}=\cap\,\{\,C\in\widehat{\mathcal{A}}\mid\Lambda_{C}^{\emptyset}\geq\Lambda(A_{D})\geq\Lambda_{\emptyset}^{C}\,\}.$
Moreover, from the proof we also conclude:
  • If $D_{0}=\emptyset$ or is located away from $\partial D^{\bullet}$, and in Assumption 12.1(ii) there is only positive definiteness near $\partial D^{\bullet}$, we only need to check the inequality $\Lambda(A_{D})\geq\Lambda_{\emptyset}^{C}$.
  • If $D_{\infty}=\emptyset$ or is located away from $\partial D^{\bullet}$, and in Assumption 12.1(ii) there is only negative definiteness near $\partial D^{\bullet}$, we only need to check the inequality $\Lambda_{C}^{\emptyset}\geq\Lambda(A_{D})$.
We also state the following two results for the inner approach of the monotonicity method, to either determine perfectly insulating or perfectly conducting inclusions. We do not give proofs of these results, as they are relatively short adaptations of the proofs of [7, theorems 3.4 and 3.5], using the tools introduced in the present paper for the extreme anisotropic conductivities.
To shorten the notation, we introduce the test-operators:
$\Lambda_{c,B}$ $=\Lambda(A_{0}+c\chi_{B}I),$
$\textup{D}\mkern-1.5mu\Lambda_{c,B}$ $=\textup{D}\mkern-1.5mu\Lambda(A_{0};c\chi_{B}I),$
where $\chi_{B}$ is a characteristic function on the set $B$.
Theorem 12.4.
Assume $D\in\widehat{\mathcal{A}}$, $A_{0}$ satisfies the UCP, and let $A_{0}\geq cI$ in the Loewner order in $\Omega$ for some $c>0$. For any open set $B\subseteq\Omega$ we have
$B\subset D\quad$ $\text{if and only if}\quad\textup{D}\mkern-1.5mu\Lambda_{-c,B}\leq\Lambda_{D}^{\emptyset}-\Lambda(A_{0})$
$\text{if and only if}\quad\Lambda_{\emptyset}^{D}-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\Lambda_{c,B}.$
Theorem 12.5.
Assume $D\in\widehat{\mathcal{A}}$, $A_{0}$ satisfies the UCP, and let $A_{0}\geq\alpha I$ in the Loewner order in $\Omega$ for some $\alpha>0$. For any open set $B\subseteq\Omega$ we have
$\text{If $0<c<\alpha$, then}\quad B\subset D\quad$ $\text{if and only if}\quad\Lambda_{-c,B}\leq\Lambda_{D}^{\emptyset}.$
$\text{If $c>0$, then}\quad B\subset D\quad$ $\text{if and only if}\quad\Lambda_{\emptyset}^{D}\leq\Lambda_{c,B}.$
13. Extreme measurements as a limit
Throughout this section, $C_{0}$ and $C_{\infty}$ satisfy Assumption 11.1 and $C=C_{0}\cup C_{\infty}$. Moreover, AA is given by (11.1) with $A_{\textup{F}}\in\mathcal{H}_{\textup{SA}}(\Omega)$.
We define $P$ as the orthogonal projection of $\mathscr{H}_{C_{0}}^{\emptyset}$ onto $\mathscr{H}_{C_{0}}^{C_{\infty}}$, and define $P^{\perp}$ as the orthogonal projection of $\mathscr{H}_{C_{0}}^{\emptyset}$ onto $(\mathscr{H}_{C_{0}}^{C_{\infty}})^{\perp}$, both with respect to the inner product $\langle\,\cdot\,,\,\cdot\,\rangle_{A}$.
Lemma 13.1.
Let $f\in L^{2}_{\diamond}(\Gamma)$ and let $u_{f}^{A_{\textup{F}}}$ satisfy (2.2) with finite conductivity $A_{\textup{F}}$. Then the solution $u_{f}^{A}$ to (11.2) is characterized as
$u_{f}^{A}=P(u_{f}^{A_{\textup{F}}}|_{\Omega\setminus C_{0}}-w),$
for $w\in H_{\diamond}^{1}(\Omega\setminus C_{0})$ solving
[…]
$\nu\cdot(A_{\textup{F}}\nabla w)$ $=\nu\cdot(A_{\textup{F}}\nabla u_{f}^{A_{\textup{F}}})\textup{ on }\partial C_{0}$
$\nu\cdot(A_{\textup{F}}\nabla w)$ $=0\textup{ on }\partial\Omega.$
Proof.
It is evident that $\widehat{u}=u_{f}^{A_{\textup{F}}}|_{\Omega\setminus C_{0}}-w$ satisfies the corresponding conductivity equation in $\Omega\setminus C_{0}$, but with vanishing Neumann condition on $\partial C_{0}$ and Neumann condition $f$ on $\partial\Omega$. This corresponds to having the perfectly insulating inclusion in $C_{0}$, but still missing the perfectly conducting inclusion in $C_{\infty}$.
Since $P$ is self-adjoint, the weak form for $\widehat{u}$ implies
$\langle f,v\rangle=\langle\widehat{u},v\rangle_{A}=\langle\widehat{u},Pv\rangle_{A}=\langle P\widehat{u},v\rangle_{A}$
for all $v\in\mathscr{H}_{C_{0}}^{C_{\infty}}\subseteq\mathscr{H}_{C_{0}}^{\emptyset}$. Hence $u_{f}^{A}=P\widehat{u}$ is the unique solution to the weak problem (11.2). ∎
Lemma 13.2.
Let $v\in\mathscr{H}_{C_{0}}^{\emptyset}$ satisfy $\nabla\cdot(A_{\textup{F}}\nabla v)=0$ in $\Omega\setminus C$. There exists $K>0$ (independent of $v$) such that
$\lVert P^{\perp}v\rVert_{A}\leq K\lVert\nabla v\rVert_{L^{2}(C_{\infty})^{d}}.$
Proof.
We let $\widehat{v}$ be the piecewise constant function in $C_{\infty}$, which in each component $C_{i}$ of $C_{\infty}$ equals the mean value of $v|_{C_{i}}$. Moreover, let $w\in\mathscr{H}_{C_{0}}^{C_{\infty}}$ solve the Dirichlet problem
$-\nabla\cdot(A_{\textup{F}}\nabla w)$ $=0\text{ in }\Omega\setminus C$
$w$ $=v\text{ on }\partial(\Omega\setminus C_{0})$
$w$ $=\widehat{v}\text{ in }C_{\infty}^{\circ}.$
Evidently vwv-w satisfies a similar PDE problem in $\Omega\setminus C$ whose only non-trivial Dirichlet condition is on $\partial C_{\infty}$. We may now estimate vwv-w using the continuous dependence on the Dirichlet condition, the trace theorem in $C_{\infty}^{\circ}$, and the Poincaré inequality applied in each component $C_{i}$:
$\lVert v-w\rVert_{H^{1}(\Omega\setminus C_{0})}^{2}$ $=\lVert v-w\rVert_{H^{1}(\Omega\setminus C)}^{2}+\lVert v-\widehat{v}\rVert_{H^{1}(C_{\infty}^{\circ})}^{2}$
$\leq K\bigl(\lVert v-w\rVert_{H^{1/2}(\partial C_{\infty})}^{2}+\lVert v-\widehat{v}\rVert_{H^{1}(C_{\infty}^{\circ})}^{2}\bigr)$
$\leq K\lVert v-\widehat{v}\rVert_{H^{1}(C_{\infty}^{\circ})}^{2}$
$\leq K\lVert\nabla v\rVert_{L^{2}(C_{\infty})^{d}}^{2}.$
The proof is concluded using the projection theorem and norm equivalences:
$\lVert P^{\perp}v\rVert_{A}=\inf_{\widetilde{v}\in\mathscr{H}_{C_{0}}^{C_{\infty}}}\lVert v-\tilde{v}\rVert_{A}\leq\lVert v-w\rVert_{A}\leq K\lVert\nabla v\rVert_{L^{2}(C_{\infty})^{d}}.\qed$
These lemmas establish a convergence result for obtaining electric potentials and local ND maps for extreme coefficients as limits, based on a sequence of truncated coefficients. Recall Definition 11.2 for the extension operator $E$.
Theorem 13.3.
Let $\epsilon>0$ and consider the truncated conductivity coefficient
$A_{\epsilon}=\begin{dcases}A_{\textup{F}}&\text{in }\Omega\setminus C\\ \epsilon A_{\textup{F}}&\text{in }C_{0}\\ \epsilon^{-1}A_{\textup{F}}&\text{in }C_{\infty}.\end{dcases}$
There exists $K>0$ (only depending on $\Omega$, $_{0}, CC_{\infty}$, and bounds on $A_{\textup{F}}$) such that
$\lVert Eu_{f}^{A}-u_{f}^{A_{\epsilon}}\rVert_{H^{1}(\Omega)}\leq K\epsilon^{1/2}\lVert f\rVert,$
and consequently
$\lVert\Lambda(A)-\Lambda(A_{\epsilon})\rVert_{\mathscr{L}(L^{2}_{\diamond}(\Gamma))}\leq K\epsilon^{1/2}.$
Remark 13.4.
It is straightforward to modify the proof of Theorem 13.3 to obtain the same estimates in case the coefficient is only partially truncated, i.e. for
$A_{\epsilon}=\begin{dcases}A_{\textup{F}}&\text{in }\Omega\setminus C\\ \epsilon A_{\textup{F}}&\text{in }C_{0}\\ \infty I&\text{in }C_{\infty}\end{dcases}\quad\text{or}\quad A_{\epsilon}=\begin{dcases}A_{\textup{F}}&\text{in }\Omega\setminus C\\ 0I&\text{in }C_{0}\\ \epsilon^{-1}A_{\textup{F}}&\text{in }C_{\infty}.\end{dcases}$
Proof.
To ease the presentation let $u_{\epsilon}=u_{f}^{A_{\epsilon}}$ and $u=u_{f}^{A}$, and let $c>0$ such that $A_{\textup{F}}\geq cI$ in the Loewner order in $\Omega$. We estimate $u_{\epsilon}$ in $C_{0}$ and $C_{\infty}$ by using the “Cauchy inequality with $\epsilon$” (here with $\alpha>0$ taking the role of the $\epsilon$ in the inequality) combined with the trace theorem in $\Omega\setminus C$:
$c\bigl(\lVert\nabla u_{\epsilon}\rVert_{L^{2}(\Omega\setminus C)}^{2}+\epsilon\lVert\nabla u_{\epsilon}\rVert_{L^{2}(C_{0})}^{2}+\epsilon^{-1}\lVert\nabla u_{\epsilon}\rVert_{L^{2}(C_{\infty})}^{2}\bigr)$ $\leq\int_{\Omega}A_{\epsilon}\nabla u_{\epsilon}\cdot\overline{\nabla u_{\epsilon}}\,\mathrm{d}x$
$=\langle f,u_{\epsilon}\rangle$
$\leq\alpha\lVert f\rVert^{2}+\frac{1}{4\alpha}\lVert u_{\epsilon}\rVert_{H^{1/2}(\partial(\Omega\setminus C))}^{2}$
$\leq\alpha\lVert f\rVert^{2}+\frac{\widehat{K}}{4\alpha}\lVert\nabla u_{\epsilon}\rVert_{L^{2}(\Omega\setminus C)^{d}}^{2}.$
Choosing $\alpha=\frac{\widehat{K}}{4c}$ results in
$\epsilon\lVert\nabla u_{\epsilon}\rVert_{L^{2}(C_{0})^{d}}^{2}+\epsilon^{-1}\lVert\nabla u_{\epsilon}\rVert_{L^{2}(C_{\infty})^{d}}^{2}\leq\frac{\widehat{K}}{4c^{2}}\lVert f\rVert^{2},$
or rather there is a constant $K>0$ such that
$\lVert\nabla u_{\epsilon}\rVert_{L^{2}(C_{\infty})^{d}}$ $\leq K\epsilon^{1/2}\lVert f\rVert,$ (13.1)
$\lVert\nabla u_{\epsilon}\rVert_{L^{2}(C_{0})^{d}}$ $\leq K\epsilon^{-1/2}\lVert f\rVert.$ (13.2)
We will make use of these bounds momentarily. First we consider $u_{\epsilon}|_{\Omega\setminus C_{0}}\in\mathscr{H}_{C_{0}}^{\emptyset}$ and apply the self-adjointness of the projections $P$ and $P^{\perp}$, and the Cauchy–Schwarz inequality, to estimate $u-u_{\epsilon}$:
$\lVert u-u_{\epsilon}\rVert_{A}^{2}$ $=\lVert P(u-u_{\epsilon})\rVert_{A}^{2}+\lVert P^{\perp}(u-u_{\epsilon})\rVert_{A}^{2}$
$\leq\langle u-u_{\epsilon},P(u-u_{\epsilon})\rangle_{A}+\lVert P^{\perp}u_{\epsilon}\rVert_{A}\lVert u-u_{\epsilon}\rVert_{A}.$ (13.3)
A combination of Lemma 13.2 and (13.1) gives the estimate
$\lVert P^{\perp}u_{\epsilon}\rVert_{A}\lVert u-u_{\epsilon}\rVert_{A}\leq K\epsilon^{1/2}\lVert f\rVert\lVert u-u_{\epsilon}\rVert_{A}.$ (13.4)
What remains is to bound the first term in (13.3) in a similar manner.
Since $P(u-u_{\epsilon})$ has vanishing gradient in $C_{\infty}^{\circ}$, the weak form for $u_{\epsilon}$ gives
$\langle u_{\epsilon},P(u-u_{\epsilon})\rangle_{A}=\langle f,P(u-u_{\epsilon})\rangle-\epsilon\int_{C_{0}}A_{\textup{F}}\nabla u_{\epsilon}\cdot\overline{\nabla w}\,\mathrm{d}x,$
where $w$ is an $H^{1}$-extension of $P(u-u_{\epsilon})$ onto $C_{0}$. Hence from the weak form of $u$, we have
$\langle u-u_{\epsilon},P(u-u_{\epsilon})\rangle_{A}=\epsilon\int_{C_{0}}A_{\textup{F}}\nabla u_{\epsilon}\cdot\overline{\nabla w}\,\mathrm{d}x.$ (13.5)
The boundedness of the $H^{1}$-extension operator and the projection $P$, together with norm equivalences, imply
$\lVert\nabla w\rVert_{L^{2}(C_{0})^{d}}\leq\lVert w\rVert_{H^{1}(\Omega)}\leq K\lVert u-u_{\epsilon}\rVert_{A}.$
Combining this with (13.5) and (13.2) leads to the estimate
$\langle u-u_{\epsilon},P(u-u_{\epsilon})\rangle_{A}\leq K\epsilon^{1/2}\lVert f\rVert\lVert u-u_{\epsilon}\rVert_{A}.$ (13.6)
Finally, (13.3), (13.4), and (13.6), together with norm equivalences, results in
$\lVert u-u_{\epsilon}\rVert_{H^{1}(\Omega\setminus C_{0})}\leq K\epsilon^{1/2}\lVert f\rVert.$ (13.7)
We now consider the extension $Eu$ and note that $v=(Eu-u_{\epsilon})|_{C_{0}}$ is a weak solution of
[…]
$v$ $=u-u_{\epsilon}\text{ on }\partial C_{0}.$
From the continuous dependence on the Dirichlet condition, the trace theorem in $\Omega\setminus C_{0}$, and (13.7) we have
$\lVert Eu-u_{\epsilon}\rVert_{H^{1}(C_{0}^{\circ})}\leq K\lVert u-u_{\epsilon}\rVert_{H^{1/2}(\partial C_{0})}\leq K\lVert u-u_{\epsilon}\rVert_{H^{1}(\Omega\setminus C_{0})}\leq K\epsilon^{1/2}\lVert f\rVert.$
Combined with (13.7), this proves the first part of the theorem.
The estimate for the local ND maps follows from (13.7) and the trace theorem in $\Omega\setminus C_{0}$:
$\lVert\Lambda(A)-\Lambda(A_{\epsilon})\rVert_{\mathscr{L}(L_{\diamond}^{2}(\Gamma))}=\sup_{\lVert f\rVert=1}\lVert u-u_{\epsilon}\rVert\leq K\sup_{\lVert f\rVert=1}\lVert u-u_{\epsilon}\rVert_{H^{1}(\Omega\setminus C_{0})}\leq K\epsilon^{1/2}.\qed$
14. Extreme operator inequalities
Throughout this section, $C_{0}$ and $C_{\infty}$ satisfy Assumption 11.1 and $C=C_{0}\cup C_{\infty}$. We re-emphasize that $C_{0}$ and $C_{\infty}$ are allowed to be empty sets, since this will be particularly relevant when using the following operator inequalities in the proof of Theorem 12.2 in Section 16. We also let $f\in L^{2}_{\diamond}(\Gamma)$ as usual.
Proposition 14.1.
Let $A_{1},A_{2}\in\mathcal{H}_{\textup{SA}}(\Omega)$ and for $j\in\{1,2\}$ define
$\widehat{A}_{j}=\begin{dcases}A_{j}&\text{in }\Omega\setminus C\\ 0I&\text{in }C_{0}\\ \infty I&\text{in }C_{\infty}.\end{dcases}$
We denote $\Lambda_{j}=\Lambda(\widehat{A}_{j})$ and $u_{j}=u_{f}^{\widehat{A}_{j}}$. Then
$\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle$ $\geq\int_{\Omega\setminus C}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x,$ (14.1)
$\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle$ $\leq\int_{\Omega\setminus C}A_{2}A_{1}^{-1}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x.$ (14.2)
Proof.
Using the weak forms, we obtain
$\langle u_{1},u_{2}\rangle_{\widehat{A}_{1}}=\langle f,\Lambda_{2}f\rangle=\lVert u_{2}\rVert_{\widehat{A}_{2}}^{2}.$
Therefore
$0$ $\leq\lVert u_{1}-u_{2}\rVert_{\widehat{A}_{1}}^{2}$
$=\lVert u_{1}\rVert_{\widehat{A}_{1}}^{2}+\lVert u_{2}\rVert_{\widehat{A}_{1}}^{2}-2\mathop{\textup{Re}}\langle u_{1},u_{2}\rangle_{\widehat{A}_{1}}$
$=\langle f,\Lambda_{1}f\rangle-\langle f,\Lambda_{2}f\rangle+\lVert u_{2}\rVert_{\widehat{A}_{1}}^{2}-\lVert u_{2}\rVert_{\widehat{A}_{2}}^{2},$ (14.3)
which is just a different way of stating (14.1), since $u_{j}$ has vanishing gradient in $C_{\infty}^{\circ}$.
Next we prove (14.2):
$\int_{\Omega\setminus C}A_{2}A_{1}^{-1}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x$
$\hskip 42.67912pt\geq\int_{\Omega\setminus C}A_{2}A_{1}^{-1}(A_{2}-A_{1})\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x-\int_{\Omega\setminus C}\lvert A_{1}^{1/2}\nabla u_{1}-A_{1}^{-1/2}A_{2}\nabla u_{2}\rvert^{2}\,\mathrm{d}x$
$\hskip 42.67912pt=-\int_{\Omega\setminus C}A_{2}\nabla u_{2}\cdot\overline{\nabla u_{2}}\,\mathrm{d}x-\int_{\Omega\setminus C}A_{1}\nabla u_{1}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x+2\mathop{\textup{Re}}\Bigl(\int_{\Omega\setminus C}A_{2}\nabla u_{2}\cdot\overline{\nabla u_{1}}\,\mathrm{d}x\Bigr)$
$\hskip 42.67912pt=\langle f,(\Lambda_{1}-\Lambda_{2})f\rangle.\qed$
In the following two results we let $A$ be as in (11.1) with $A_{\textup{F}}\in\mathcal{H}_{\textup{SA}}(\Omega)$ and with the extreme inclusion $C_{0}$ and $C_{\infty}$. We also introduce the short-hand notation $\Lambda_{C_{0}}^{C_{\infty}}=\Lambda(A)$ for the local ND map and $u_{C_{0}}^{C_{\infty}}=u_{f}^{A}$ for the electric potential.
Proposition 14.2.
There exists $K>0$ (independent of $f$) such that
$\langle f,(\Lambda_{C_{0}}^{\emptyset}-\Lambda_{C_{0}}^{C_{\infty}})f\rangle$ $\geq\int_{C_{\infty}}A_{\textup{F}}\nabla u_{C_{0}}^{\emptyset}\cdot\overline{\nabla u_{C_{0}}^{\emptyset}}\,\mathrm{d}x$ (14.4)
$\langle f,(\Lambda_{C_{0}}^{\emptyset}-\Lambda_{C_{0}}^{C_{\infty}})f\rangle$ $\leq K\int_{C_{\infty}}\lvert\nabla u_{C_{0}}^{\emptyset}\rvert^{2}\,\mathrm{d}x.$ (14.5)
Proof.
We start by proving (14.4). We define a truncated coefficient (in $C_{\infty}$), for $\epsilon>0$:
$A_{\epsilon}=\begin{dcases}A_{\textup{F}}&\text{in }\Omega\setminus C\\ 0I&\text{in }C_{0}\\ \epsilon^{-1}A_{\textup{F}}&\text{in }C_{\infty}.\end{dcases}$
Using (14.2) from Proposition 14.1 (here with $\Omega\setminus C$ replaced by $\Omega\setminus C_{0}$), we have
$\langle f,\bigl(\Lambda_{C_{0}}^{\emptyset}-\Lambda(A_{\epsilon})\bigr)f\rangle\geq(1-\epsilon)\int_{C_{\infty}}A_{\textup{F}}\nabla u_{C_{0}}^{\emptyset}\cdot\overline{\nabla u_{C_{0}}^{\emptyset}}\,\mathrm{d}x.$
We arrive at (14.4) by letting $\epsilon\to 0$ due to Theorem 13.3 (Remark 13.4).
Next we consider (14.5). Note that the proof of Lemma 13.1 implies that $u_{C_{0}}^{C_{\infty}}=Pu_{C_{0}}^{\emptyset}$, and therefore $P^{\perp}u_{C_{0}}^{\emptyset}=u_{C_{0}}^{\emptyset}-u_{C_{0}}^{C_{\infty}}$. Inserting this into the weak form for $u_{C_{0}}^{\emptyset}$ gives
$\langle f,(\Lambda_{C_{0}}^{\emptyset}-\Lambda_{C_{0}}^{C_{\infty}})f\rangle=\langle f,P^{\perp}u_{C_{0}}^{\emptyset}\rangle=\langle u_{C_{0}}^{\emptyset},P^{\perp}u_{C_{0}}^{\emptyset}\rangle_{A}=\lVert P^{\perp}u_{C_{0}}^{\emptyset}\rVert_{A}^{2}.$
Hence, the result follows by applying Lemma 13.2. ∎
Proposition 14.3.
We have the bounds
$\langle f,(\Lambda_{C_{0}}^{C_{\infty}}-\Lambda_{\emptyset}^{C_{\infty}})f\rangle$ $\geq\int_{C_{0}}A_{\textup{F}}\nabla u_{\emptyset}^{C_{\infty}}\cdot\overline{\nabla u_{\emptyset}^{C_{\infty}}}\,\mathrm{d}x$ (14.6)
$\langle f,(\Lambda_{C_{0}}^{C_{\infty}}-\Lambda_{\emptyset}^{C_{\infty}})f\rangle$ $\leq\int_{C_{0}}A_{\textup{F}}\nabla Eu_{C_{0}}^{C_{\infty}}\cdot\overline{\nabla Eu_{C_{0}}^{C_{\infty}}}\,\mathrm{d}x.$ (14.7)
Proof.
Once again we start by introducing a truncated coefficient (this time in $C_{0}$):
$A_{\epsilon}=\begin{dcases}A_{\textup{F}}&\text{in }\Omega\setminus C\\ \epsilon A_{\textup{F}}&\text{in }C_{0}\\ \infty I&\text{in }C_{\infty}.\end{dcases}$
Using (14.1) from Proposition 14.1 (here with $\Omega\setminus C$ replaced by $\Omega\setminus C_{\infty}$), we have
$\langle f,\bigl(\Lambda(A_{\epsilon})-\Lambda_{\emptyset}^{C_{\infty}}\bigr)f\rangle\geq(1-\epsilon)\int_{C_{0}}A_{\textup{F}}\nabla u_{\emptyset}^{C_{\infty}}\cdot\overline{\nabla u_{\emptyset}^{C_{\infty}}}\,\mathrm{d}x.$
We arrive at (14.6) by letting $\epsilon\to 0$ due to Theorem 13.3 (Remark 13.4).
Next we consider (14.7). Let $\widehat{A}$ be the coefficient associated with $\Lambda_{\emptyset}^{C_{\infty}}$, recall the extension operator $E$ (Definition 11.2), and note that
$\langle u_{\emptyset}^{C_{\infty}},Eu_{C_{0}}^{C_{\infty}}\rangle_{\widehat{A}}=\langle f,\Lambda_{C_{0}}^{C_{\infty}}f\rangle=\lVert Eu_{C_{0}}^{C_{\infty}}\rVert_{\widehat{A}}^{2}-\int_{C_{0}}A_{\textup{F}}\nabla Eu_{C_{0}}^{C_{\infty}}\cdot\overline{\nabla Eu_{C_{0}}^{C_{\infty}}}\,\mathrm{d}x.$
The inequality (14.7) is therefore given by the following computation:
$0$ $\leq\lVert Eu_{C_{0}}^{C_{\infty}}-u_{\emptyset}^{C_{\infty}}\rVert_{\widehat{A}}^{2}$
$=\lVert Eu_{C_{0}}^{C_{\infty}}\rVert_{\widehat{A}}^{2}+\lVert u_{\emptyset}^{C_{\infty}}\rVert_{\widehat{A}}^{2}-2\mathop{\textup{Re}}\langle u_{\emptyset}^{C_{\infty}},Eu_{C_{0}}^{C_{\infty}}\rangle_{\widehat{A}}$
$=\langle f,(\Lambda_{\emptyset}^{C_{\infty}}-\Lambda_{C_{0}}^{C_{\infty}})f\rangle+\int_{C_{0}}A_{\textup{F}}\nabla Eu_{C_{0}}^{C_{\infty}}\cdot\overline{\nabla Eu_{C_{0}}^{C_{\infty}}}\,\mathrm{d}x.\qed$
15. Localized potentials for extreme coefficients
Throughout this section, $C_{0}$ and $C_{\infty}$ satisfy Assumption 11.1 and $C=C_{0}\cup C_{\infty}$. Moreover, AA is given by (11.1) with $A_{\textup{F}}\in\mathcal{H}_{\textup{SA}}(\Omega)$.
We have previously introduced localized potentials for anisotropic coefficients in Theorem 6.3, and how one can transfer the localization between different coefficients in Propostion 6.4, provided the coefficients only differ in the set where the power tends to zero. Below we show that such localization can also be transferred to the case of extreme inclusions, again provided that the extreme inclusions are located in the set where the power tends to zero.
Proposition 15.1.
Let $B$ and $U$ be as in Theorem 6.3, and assume for $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ that $u_{j}=u_{f_{j}}^{A_{\textup{F}}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=0.$
If $C\subset\Omega\setminus\overline{U}$ then $\widehat{u}_{j}=Eu_{f_{j}}^{A}$ also satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x=0.$
Proof.
Let $v_{j}=u_{j}|_{\Omega\setminus C_{0}}-w_{j}$ for $w_{j}$ solving the auxiliary PDE problem from Lemma 13.1 (with $f$ replaced by $f_{j}$), and consider $\widehat{v}_{j}=u_{f_{j}}^{A}=Pv_{j}$.
Let $V$ be an open set satisfying $C_{0}\subset\Omega\setminus\overline{V}\subseteq\Omega\setminus\overline{U}$, and such that $\partial(\Omega\setminus\overline{V})$ is Lipschitz regular. By continuous dependence of $w_{j}$ on its Neumann condition and from the weak form of $u_{j}|_{\Omega\setminus(V\cup C_{0})}$, we have the estimate
$\lVert w_{j}\rVert_{H^{1}(\Omega\setminus C_{0})}$ $\leq K\lVert\nu\cdot(A_{\textup{F}}\nabla u_{j})\rVert_{H^{-1/2}(\partial C_{0})}$
$\leq K\lVert\nu\cdot(A_{\textup{F}}\nabla u_{j})\rVert_{H^{-1/2}(\partial(\Omega\setminus(V\cup C_{0})))}$
$\leq K\lVert\nabla u_{j}\rVert_{L^{2}(\Omega\setminus(U\cup C_{0}))^{d}}\to 0\text{ for }j\to\infty.$
Consequently, from the limits of $u_{j}$ and $w_{j}$, we have
$\lim_{j\to\infty}\int_{B}\lvert\nabla v_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus(U\cup C_{0})}\lvert\nabla v_{j}\rvert^{2}\,\mathrm{d}x=0,$ (15.1)
and from Lemma 13.2 we have
$\lVert P^{\perp}v_{j}\rVert_{A}\leq K\lVert\nabla v_{j}\rVert_{L^{2}(C_{\infty})^{d}}\to 0\text{ for }j\to\infty.$ (15.2)
Since $\widehat{v}_{j}=v_{j}-P^{\perp}v_{j}$, then a combination of (15.1) and (15.2) gives
$\lim_{j\to\infty}\int_{B}\lvert\nabla\widehat{v}_{j}\rvert^{2}\,\mathrm{d}x=\infty\qquad\text{and}\qquad\lim_{j\to\infty}\int_{\Omega\setminus(U\cup C_{0})}\lvert\nabla\widehat{v}_{j}\rvert^{2}\,\mathrm{d}x=0.$
We have $\widehat{u}_{j}=E\widehat{v}_{j}$, and have thus proven the limiting behavior in $\Omega\setminus C_{0}$. What remains is to investigate $\widehat{u}_{j}|_{C_{0}}$.
In the following we may assume that $\Omega\setminus(\overline{V}\cup C_{0})$ is connected, otherwise a similar estimate can be achieved by working in each component separately that touches $C_{0}$. Let $c_{j}$ denote the mean of $\widehat{u}_{j}$ in $\Omega\setminus(\overline{V}\cup C_{0})$, then from the trace theorem and Poincaré inequality we have
$\lVert\widehat{u}_{j}-c_{j}\rVert_{H^{1/2}(\partial C_{0})}$ $\leq K\lVert\widehat{u}_{j}-c_{j}\rVert_{H^{1}(\Omega\setminus(\overline{V}\cup C_{0}))}$
$\leq K\lVert\nabla\widehat{u}_{j}\rVert_{L^{2}(\Omega\setminus(V\cup C_{0}))^{d}}\to 0\text{ for }j\to\infty.$
However, as $\widehat{u}_{j}-c_{j}$ solves the same PDE as $\widehat{u}_{j}$ in $C_{0}$, we conclude from the continuous dependence on the Dirichlet condition that
$\lVert\nabla\widehat{u}_{j}\rVert_{L^{2}(C_{0})^{d}}=\lVert\nabla(\widehat{u}_{j}-c_{j})\rVert_{L^{2}(C_{0})^{d}}\leq K\lVert\widehat{u}_{j}-c_{j}\rVert_{H^{1/2}(\partial C_{0})}\to 0\text{ for }j\to\infty.\qed$
16. Proof of Theorem 12.2
Recall that $A_{0}\in\mathcal{H}_{\textup{SA}}(\Omega)$ and the unknown conductivity $A_{D}$ is of the form
[…]
where $D_{0}$ and $D_{\infty}$ satisfy Assumption 11.1 (in place of $C_{0}$ and $C_{\infty}$), and with the finite part $A_{\textup{F}}\in\mathcal{H}_{\textup{SA}}(\Omega)$. The inclusions are
$D=\mathop{\textup{supp}}(A_{D}-A_{0}),$
and recall that
$D_{\textup{F}}=D\setminus(D_{0}\cup D_{\infty})$
is the part of the domain occupied by non-extreme deviations from $A_{0}$.
Proof of $\boldsymbol{D\subseteq C\Rightarrow\Lambda_{C}^{\emptyset}\geq\Lambda(A_{D})\geq\Lambda_{\emptyset}^{C}}$
We truncate $A_{D}$ for an $\epsilon>0$ as
$A_{\epsilon}=\begin{dcases}A_{\textup{F}}&\text{in }\Omega\setminus(D_{0}\cup D_{\infty})\\ \epsilon A_{\textup{F}}&\text{in }D_{0}\\ \epsilon^{-1}A_{\textup{F}}&\text{in }D_{\infty},\end{dcases}$
and similarly truncate the test-coefficients
$\widehat{A}_{\epsilon}=\begin{dcases}A_{\textup{F}}&\text{in }\Omega\setminus C\\ \epsilon A_{\textup{F}}&\text{in }C\end{dcases}\qquad\text{and}\qquad\widetilde{A}_{\epsilon}=\begin{dcases}A_{\textup{F}}&\text{in }\Omega\setminus C\\ \epsilon^{-1}A_{\textup{F}}&\text{in }C.\end{dcases}$
For $D\subseteq C$ and $\epsilon\in(0,1]$ we have $\widehat{A}_{\epsilon}\leq A_{\epsilon}\leq\widetilde{A}_{\epsilon}$, which by (5.5) implies $\Lambda(\widehat{A}_{\epsilon})\geq\Lambda(A_{\epsilon})\geq\Lambda(\widetilde{A}_{\epsilon})$. Letting $\epsilon\to 0$, the operator norm convergence from Theorem 13.3 gives
$\Lambda_{C}^{\emptyset}\geq\Lambda(A_{D})\geq\Lambda_{\emptyset}^{C}.$
Proof of $\boldsymbol{\Lambda_{C}^{\emptyset}\geq\Lambda(A_{D})\geq\Lambda_{\emptyset}^{C}\Rightarrow D\subseteq C}$
Now assume that $D\not\subseteq C$, i.e. that $D^{\bullet}\not\subseteq C$. Since both sets are closures of opens sets and have connected complements, there is a relatively open set that connects $D^{\bullet}\setminus C$ with $\Gamma$ that we may use for localization. Specifically, there exist a relatively open connected set $U\subset\overline{\Omega}\setminus C$ that intersects $\Gamma$ and an open ball $B\subset U\cap D$. Moreover, Assumption 12.1(ii) ensures that we can pick $B$ and $U$ such that we arrive at one of four cases:
  • Case 1: $U\cap D\subseteq D_{\infty}$.
  • Case 2: $U\cap D\subseteq D_{\textup{F}}$, and $A_{\textup{F}}-A_{0}$ is positive semidefinite in $U$ and uniformly positive definite in $B$.
  • Case 3: $U\cap D\subseteq D_{0}$.
  • Case 4: $U\cap D\subseteq D_{\textup{F}}$, and $A_{\textup{F}}-A_{0}$ is negative semidefinite in $U$ and uniformly negative definite in $B$.
In the first two cases we will prove that $\Lambda(A_{D})\not\geq\Lambda_{\emptyset}^{C}$, and in the last two case we will prove that $\Lambda_{C}^{\emptyset}\not\geq\Lambda(A_{D})$. Together the four cases prove the contrapositive formulation of the assertion
$\Lambda_{C}^{\emptyset}\geq\Lambda(A_{D})\geq\Lambda_{\emptyset}^{C}\quad\text{\emph{implies}}\quad D\subseteq C.$
In the proofs of the four cases we will have several comparisons of coefficients, so we use the convention that
$A_{\textup{F}}=A_{0}\text{ in }D_{0}\cup D_{\infty}$
to simplify the presentation (since the values of $A_{\textup{F}}$ in $D_{0}$ and $D_{\infty}$ are not specified from the definition of $A_{D}$). We will also make use of the bounds
$\alpha I\leq A_{0}\leq\beta I\quad\text{and}\quad\alpha I\leq A_{\textup{F}}\leq\beta I$
in the Loewner order in $\Omega$, for suitable $\alpha,\beta>0$.
Case 1
We define two auxiliary coefficients:
$A_{1}=A_{\textup{F}}\qquad\text{and}\qquad A_{2}=\begin{dcases}A_{\textup{F}}&\text{in }\Omega\setminus D_{0}\\ 0I&\text{in }D_{0}.\end{dcases}$
We shorten the notation to $\Lambda=\Lambda(A_{D})$ and $\Lambda_{k}=\Lambda(A_{k})$ for $k\in\{0,1,2\}$. From Theorem 6.3 and Propositions 6.4 and 15.1, we may pick a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$, such that $u_{j}=u_{f_{j}}^{A_{0}}$, $\widehat{u}_{j}=u_{f_{j}}^{A_{1}}$, and $\widetilde{u}_{j}=Eu_{f_{j}}^{A_{2}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{B}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{B}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x=\infty$
and
$\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x=0.$
Next, we write
$\Lambda-\Lambda_{\emptyset}^{C}=(\Lambda-\Lambda_{2})+(\Lambda_{2}-\Lambda_{1})+(\Lambda_{1}-\Lambda_{0})+(\Lambda_{0}-\Lambda_{\emptyset}^{C}),$ (16.1)
and estimate each of the differences of the local ND maps. From Propositions 14.114.3 we have:
$\langle f_{j},(\Lambda-\Lambda_{2})f_{j}\rangle$ $\leq-\int_{D_{\infty}}A_{0}\nabla\widetilde{u}_{j}\cdot\overline{\nabla\widetilde{u}_{j}}\,\mathrm{d}x\leq-\alpha\int_{B}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x,$
$\langle f_{j},(\Lambda_{2}-\Lambda_{1})f_{j}\rangle$ $\leq\int_{D_{0}}A_{0}\nabla\widetilde{u}_{j}\cdot\overline{\nabla\widetilde{u}_{j}}\,\mathrm{d}x\leq\beta\int_{\Omega\setminus U}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x,$
$\langle f_{j},(\Lambda_{1}-\Lambda_{0})f_{j}\rangle$ $\leq\int_{D_{\textup{F}}}(A_{0}-A_{\textup{F}})\nabla\widehat{u}_{j}\cdot\overline{\nabla\widehat{u}_{j}}\,\mathrm{d}x\leq\lVert A_{0}-A_{\textup{F}}\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x,$
$\langle f_{j},(\Lambda_{0}-\Lambda_{\emptyset}^{C})f_{j}\rangle$ $\leq K\int_{C}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x\leq K\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x.$
Hence, (16.1) gives
$\lim_{j\to\infty}\langle f_{j},\bigl(\Lambda(A_{D})-\Lambda_{\emptyset}^{C}\bigr)f_{j}\rangle=-\infty,$
so we conclude that $\Lambda(A_{D})\not\geq\Lambda_{\emptyset}^{C}$.
Case 2
We define two auxiliary coefficients:
$A_{1}=\begin{dcases}A_{0}&\text{in }\Omega\setminus D_{0}\\ 0I&\text{in }D_{0}\end{dcases}\qquad\text{and}\qquad A_{2}=\begin{dcases}A_{0}&\text{in }\Omega\setminus(D_{0}\cup D_{\infty})\\ 0I&\text{in }D_{0}\\ \infty I&\text{in }D_{\infty}.\end{dcases}$
We shorten the notation to $\Lambda=\Lambda(A_{D})$ and $\Lambda_{k}=\Lambda(A_{k})$ for $k\in\{0,1,2\}$. From Theorem 6.3 and Proposition 15.1, we may pick a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$, such that $u_{j}=u_{f_{j}}^{A_{0}}$, $\widehat{u}_{j}=Eu_{f_{j}}^{A_{1}}$, and $\widetilde{u}_{j}=Eu_{f_{j}}^{A_{2}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{B}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{B}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x=\infty$
and
$\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x=0.$
Next, we write
$\Lambda-\Lambda_{\emptyset}^{C}=(\Lambda-\Lambda_{2})+(\Lambda_{2}-\Lambda_{1})+(\Lambda_{1}-\Lambda_{0})+(\Lambda_{0}-\Lambda_{\emptyset}^{C}),$ (16.2)
and estimate each of the differences of the local ND maps. Recall that $A_{0}-A_{\textup{F}}$ is negative semidefinite in $U$ and there exists a $c>0$ such that $A_{0}-A_{\textup{F}}\leq-cI$ in $B$. From Propositions 14.114.3 and Proposition 5.6 we have:
$\langle f_{j},(\Lambda-\Lambda_{2})f_{j}\rangle$ $\leq\int_{D_{\textup{F}}}A_{0}A_{\textup{F}}^{-1}(A_{0}-A_{\textup{F}})\nabla\widetilde{u}_{j}\cdot\overline{\nabla\widetilde{u}_{j}}\,\mathrm{d}x$
$\leq\lVert A_{0}A_{\textup{F}}^{-1}(A_{0}-A_{\textup{F}})\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x-c(\tfrac{\alpha}{\beta})^{2}\int_{B}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x,$
$\langle f_{j},(\Lambda_{2}-\Lambda_{1})f_{j}\rangle$ $\leq-\int_{D_{\infty}}A_{0}\nabla\widehat{u}_{j}\cdot\overline{\nabla\widehat{u}_{j}}\,\mathrm{d}x\leq 0,$
$\langle f_{j},(\Lambda_{1}-\Lambda_{0})f_{j}\rangle$ $\leq\int_{D_{0}}A_{0}\nabla\widehat{u}_{j}\cdot\overline{\nabla\widehat{u}_{j}}\,\mathrm{d}x\leq\beta\int_{\Omega\setminus U}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x,$
$\langle f_{j},(\Lambda_{0}-\Lambda_{\emptyset}^{C})f_{j}\rangle$ $\leq K\int_{C}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x\leq K\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x.$
Hence, (16.2) gives
$\lim_{j\to\infty}\langle f_{j},\bigl(\Lambda(A_{D})-\Lambda_{\emptyset}^{C}\bigr)f_{j}\rangle=-\infty,$
so we conclude that $\Lambda(A_{D})\not\geq\Lambda_{\emptyset}^{C}$.
Case 3
We define two auxiliary coefficients:
$A_{1}=A_{\textup{F}}\qquad\text{and}\qquad A_{2}=\begin{dcases}A_{\textup{F}}&\text{in }\Omega\setminus D_{\infty}\\ \infty I&\text{in }D_{\infty},\end{dcases}$
and denote the coefficient associated with the test-operator $\Lambda_{C}^{\emptyset}$ as $A_{C}$.
We shorten the notation to $\Lambda=\Lambda(A_{D})$ and $\Lambda_{k}=\Lambda(A_{k})$ for $k\in\{0,1,2\}$. From Theorem 6.3 and Propositions 6.4 and 15.1, we may pick a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$, such that $u_{j}=u_{f_{j}}^{A_{0}}$, $\widehat{u}_{j}=u_{f_{j}}^{A_{1}}$, $\widetilde{u}_{j}=u_{f_{j}}^{A_{2}}$, and $\breve{u}_{j}=Eu_{f_{j}}^{A_{C}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{B}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{B}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{B}\lvert\nabla\breve{u}_{j}\rvert^{2}\,\mathrm{d}x=\infty$
and
$\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x=\int_{\Omega\setminus U}\lvert\nabla\breve{u}_{j}\rvert^{2}\,\mathrm{d}x=0.$
Next, we write
$\Lambda_{C}^{\emptyset}-\Lambda=(\Lambda_{C}^{\emptyset}-\Lambda_{0})+(\Lambda_{0}-\Lambda_{1})+(\Lambda_{1}-\Lambda_{2})+(\Lambda_{2}-\Lambda),$ (16.3)
and estimate each of the differences of the local ND maps. From Propositions 14.114.3 we have:
$\langle f_{j},(\Lambda_{C}^{\emptyset}-\Lambda_{0})f_{j}\rangle$ $\leq\int_{C}A_{0}\nabla\breve{u}_{j}\cdot\overline{\nabla\breve{u}_{j}}\,\mathrm{d}x\leq\beta\int_{\Omega\setminus U}\lvert\nabla\breve{u}_{j}\rvert^{2}\,\mathrm{d}x,$
$\langle f_{j},(\Lambda_{0}-\Lambda_{1})f_{j}\rangle$ $\leq\int_{D_{\textup{F}}}(A_{\textup{F}}-A_{0})\nabla u_{j}\cdot\overline{\nabla u_{j}}\,\mathrm{d}x\leq\lVert A_{\textup{F}}-A_{0}\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x,$
$\langle f_{j},(\Lambda_{1}-\Lambda_{2})f_{j}\rangle$ $\leq K\int_{D_{\infty}}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x\leq K\int_{\Omega\setminus U}\lvert\nabla\widehat{u}_{j}\rvert^{2}\,\mathrm{d}x,$
$\langle f_{j},(\Lambda_{2}-\Lambda)f_{j}\rangle$ $\leq-\int_{D_{0}}A_{0}\nabla\widetilde{u}_{j}\cdot\overline{\nabla\widetilde{u}_{j}}\,\mathrm{d}x\leq-\alpha\int_{B}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x.$
Hence, (16.3) gives
$\lim_{j\to\infty}\langle f_{j},\bigl(\Lambda_{C}^{\emptyset}-\Lambda(A_{D})\bigr)f_{j}\rangle=-\infty,$
so we conclude that $\Lambda_{C}^{\emptyset}\not\geq\Lambda(A_{D})$.
Case 4
We define two auxiliary coefficients:
$A_{1}=\begin{dcases}A_{0}&\text{in }\Omega\setminus D_{\infty}\\ \infty I&\text{in }D_{\infty}\end{dcases}\qquad\text{and}\qquad A_{2}=\begin{dcases}A_{0}&\text{in }\Omega\setminus(D_{0}\cup D_{\infty})\\ 0I&\text{in }D_{0}\\ \infty I&\text{in }D_{\infty},\end{dcases}$
and denote the coefficient associated with the test-operator $\Lambda_{C}^{\emptyset}$ as $A_{C}$.
We shorten the notation to $\Lambda=\Lambda(A_{D})$ and $\Lambda_{k}=\Lambda(A_{k})$ for $k\in\{0,1,2\}$. From Theorem 6.3 and Proposition 15.1, we may pick a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$, such that $u_{j}=u_{f_{j}}^{A_{0}}$, $\widehat{u}_{j}=u_{f_{j}}^{A_{1}}$, $\widetilde{u}_{j}=Eu_{f_{j}}^{A_{2}}$, and $\breve{u}_{j}=Eu_{f_{j}}^{A_{C}}$ satisfy
$\lim_{j\to\infty}\int_{B}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{B}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{B}\lvert\nabla\breve{u}_{j}\rvert^{2}\,\mathrm{d}x=\infty$
and
$\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x=\lim_{j\to\infty}\int_{\Omega\setminus U}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x=\int_{\Omega\setminus U}\lvert\nabla\breve{u}_{j}\rvert^{2}\,\mathrm{d}x=0.$
Next, we write
$\Lambda_{C}^{\emptyset}-\Lambda=(\Lambda_{C}^{\emptyset}-\Lambda_{0})+(\Lambda_{0}-\Lambda_{1})+(\Lambda_{1}-\Lambda_{2})+(\Lambda_{2}-\Lambda),$ (16.4)
and estimate each of the differences of the local ND maps. Recall that $A_{\textup{F}}-A_{0}$ is negative semidefinite in $U$ and there exists a $c>0$ such that $A_{\textup{F}}-A_{0}\leq-cI$ in $B$. From Propositions 14.114.3 we have:
$\langle f_{j},(\Lambda_{C}^{\emptyset}-\Lambda_{0})f_{j}\rangle$ $\leq\int_{C}A_{0}\nabla\breve{u}_{j}\cdot\overline{\nabla\breve{u}_{j}}\,\mathrm{d}x\leq\beta\int_{\Omega\setminus U}\lvert\nabla\breve{u}_{j}\rvert^{2}\,\mathrm{d}x,$
$\langle f_{j},(\Lambda_{0}-\Lambda_{1})f_{j}\rangle$ $\leq K\int_{D_{\infty}}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x\leq K\int_{\Omega\setminus U}\lvert\nabla u_{j}\rvert^{2}\,\mathrm{d}x,$
$\langle f_{j},(\Lambda_{1}-\Lambda_{2})f_{j}\rangle$ $\leq-\int_{D_{0}}A_{0}\nabla\widehat{u}_{j}\cdot\overline{\nabla\widehat{u}_{j}}\,\mathrm{d}x\leq 0,$
$\langle f_{j},(\Lambda_{2}-\Lambda)f_{j}\rangle$ $\leq\int_{D_{\textup{F}}}(A_{\textup{F}}-A_{0})\nabla\widetilde{u}_{j}\cdot\overline{\nabla\widetilde{u}_{j}}\,\mathrm{d}x$
$\leq\lVert A_{\textup{F}}-A_{0}\rVert_{*}\int_{\Omega\setminus U}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x-c\int_{B}\lvert\nabla\widetilde{u}_{j}\rvert^{2}\,\mathrm{d}x.$
Hence, (16.4) gives
$\lim_{j\to\infty}\langle f_{j},\bigl(\Lambda_{C}^{\emptyset}-\Lambda(A_{D})\bigr)f_{j}\rangle=-\infty,$
so we conclude that $\Lambda_{C}^{\emptyset}\not\geq\Lambda(A_{D})$.
Acknowledgements
The authors are supported by grant 10.46540/3120-00003B from Independent Research Fund Denmark.